### v1

# Singular value asymptotics on compact smooth Riemannian manifolds

Abstract.
Let $(X,G)$ be a $d$-dimensional compact smooth Riemannian manifold equipped with Laplace-Beltrami operator $\Delta_{G}$, and let $\Pi_{X}$ be the $C^{\ast}$-algebra obtained by locally transferring the $C^{\ast}$-algebra generated by multiplication operators and Riesz transforms on $\mathbb{R}^{d}$. Denote ${\rm sym}_{X}$ the principal symbol mapping of $\Pi_{X}$. For any $S\in\Pi_{X}$, we prove that, in the framework of $C^{\ast}$-algebra,
$\lim_{t\rightarrow\infty}t^{\frac{1}{p}}\mu(t,S(1+\Delta_{G})^{-\frac{d}{2p}})=c_{d,p}\Big\|{\rm sym}_{X}(S)\Big\|_{L_{p}(T^{\ast}X,e^{-q_{G}}d\lambda)},\quad 0<p<\infty,$
where $c_{d,p}$ is a constant, $e^{-q_{G}}$ is the canonical weight on $X$, and $d\lambda$ is the Liouville measure on the cotangent bundle $T^{\ast}X$.
2020 Mathematics Subject Classification: 58C40, 46L51, 47G30
Key words: singular value function, compact manifold, Weyl’s law, principal symbol
1. Introduction
In this paper, we establish singular value asymptotics on a $C^{\ast}$-algebra over compact Riemannian manifold. On $\mathbb{R}^{d}$, recall that $\nabla=(D_{1},\cdots,D_{d})$ and $\Delta=\sum_{k=1}^{d}D_{k}^{2}$, where $D_{k}=\frac{1}{i}\partial_{k}$. The $C^{\ast}$-subalgebra $\Pi_{\mathbb{R}^{d}}$ of $B(L_{2}(\mathbb{R}^{d}))$ (the set of bounded linear operators), introduced in [37], is generated by the algebras
$\{M_{f}:f\in\mathbb{C}+C_{0}(\mathbb{R}^{d})\},\quad\{g(\frac{\nabla}{\sqrt{\Delta}}):g\in C(\mathbb{S}^{d-1})\},$
where $M_{f}$ is defined by $M_{f}\phi=f\phi$ for $\phi\in L_{1,\text{loc}}(\mathbb{R}^{d})$. As shown in [37, 26], there exists a $\ast$-homomorphism
${\rm sym}:\Pi_{\mathbb{R}^{d}}\rightarrow C(\mathbb{S}^{d-1},\mathbb{C}+C_{0}(\mathbb{R}^{d}))$
such that, for all $f\in\mathbb{C}+C_{0}(\mathbb{R}^{d})$ and $g\in C(\mathbb{S}^{d-1})$,
${\rm sym}(M_{f})=f\otimes 1\quad\mbox{and}\quad{\rm sym}(g(\frac{\nabla}{\sqrt{\Delta}}))=1\otimes g.$
This *-homomorphism, referred to as a principal symbol mapping, generalizes the classical notion of the principal symbol for pseudodifferential operators on $\mathbb{R}^{d}$.
An operator $S\in B(L_{2}(\mathbb{R}^{d}))$ is compactly supported if there exists $\phi\in C_{c}^{\infty}(\mathbb{R}^{d})$ such that $S=M_{\phi}SM_{\phi}$. For a compact operator $T$ on Hilbert space $H$, let $t\mapsto\mu(t,T)$ be the singular value function (see Subsection 2.4 for detail). The following singular value asymptotics on $\mathbb{R}^{d}$ was proved in[13, Theorem 6.1] (earlier in [12] with the case $p=d$).
Theorem A.
For $0<p<\infty$, if $S\in\Pi_{\mathbb{R}^{d}}$ is compactly supported, then
$\lim_{t\rightarrow\infty}t^{\frac{1}{p}}\mu\Big(t,S(1+\Delta)^{-\frac{d}{2p}}\Big)=c_{d}^{\frac{1}{p}}\|{\rm sym}(S)\|_{L_{p}(\mathbb{R}^{d}\times\mathbb{S}^{d-1})}$
where $c_{d}=d^{-\frac{1}{d}}(2\pi)^{-1}$.
The Theorem A is closely connected to the Connes’ integral established in [24, Theorem 5.3.3](also [37, Theorem 1.5]), and plays a key role in studying the singular value asymptotics of quantum derivatives [12, 36, 40] and spectral estimates [13, 11].
Let $(X,G)$ be a compact smooth Riemannian manifold with Riemannian metric $G$ and an atlas $\{(\mathcal{U}_{i},h_{i})\}_{i\in\mathbb{I}}$, and let ${\rm vol}_{G}$ be the volume form on $X$ induced by the metric $G$. On the compact manifold $X$, the $C^{\ast}$-algebra $\Pi_{X}$ introduced in [22] corresponds to $\Pi_{\mathbb{R}^{d}}$ in local coordinates.
Definition 1.1.
Let $(X,G)$ be a $d$-dimensional compact smooth Riemannian manifold with Riemannian metric $G$. For $S\in B(L_{2}(X,{\rm vol}_{G}))$, we say $S\in\Pi_{X}$ if
  1. (1)
    For each $i\in\mathbb{I}$ and each $\phi\in C_{c}(\mathcal{U}_{i})$, the operator $M_{\phi}SM_{\phi}$, when transferred to an operator on $L_{2}(\mathbb{R}^{d})$, belongs to $\Pi_{\mathbb{R}^{d}}$;
  2. (2)
    For each $\psi\in C(X)$, the commutator $M_{\psi},S]$ is compact.
The $C^{\ast}$-algebra $\Pi_{\mathbb{R}^{d}}$ is strictly broader than the class of classical compactly supported pseudo-differential operators of order 0 on $\mathbb{R}^{d}$ [21]. By the definition, the algebra $\Pi_{X}$ contains the class of classical compactly supported pseudo-differential operators of order 0 in each chart.
Let $S^{\ast}X$ denote the cosphere bundle of $X$. By [22, Theorem 1.4], the algebra $\Pi_{X}$ is a $C^{\ast}$-algebra, and there exists a surjective $\ast$-homomorphism
${\rm sym}_{X}:\Pi_{X}\rightarrow C(S^{*}X)$
such that
${\rm ker}({\rm sym}_{X})=\mathcal{K}(L_{2}(X,{\rm vol}_{G})).$
Here, $\mathcal{K}(L_{2}(X,{\rm vol}_{G}))$ is the set of compact operators on Hilbert space $L_{2}(X,{\rm vol}_{G})$. We prove an analogue of Theorem A on compact Riemannian manifolds.
Theorem 1.2.
Let $(X,G)$ be a compact smooth Riemannian manifold with Riemannian metric $G$. For $0<p<\infty$, if $S\in\Pi_{X}$, then
$\lim_{t\rightarrow\infty}t^{\frac{1}{p}}\mu(t,S(1+\Delta_{G})^{-\frac{d}{2p}})=(2\pi\sqrt[d]{d})^{-\frac{1}{p}}\Big\|{\rm sym}_{X}(S)\Big\|_{L_{p}(T^{\ast}X,e^{-q_{G}}d\lambda)}$
where $\Delta_{G}$ and $e^{-q_{G}}$ are the Laplace–Beltrami operator and canonical weight on $X$ separately, and $d\lambda$ is the Liouville measure on the cotangent bundle $T^{\ast}X$.
For the precise definitions of the Laplace–Beltrami operator $\Delta_{G}$, the weight $e^{-q_{G}}$, and the Liouville measure $d\lambda$, we refer to Subsection 2.1.
The Theorem 1.2 covers the results on closed Riemannian manifolds [1, 14] and the classical tori [36] (this result is different with that in [1, 14]), and has a direct link with Connes’ integral in [26, 22] on compact Riemannian manifolds. Let us give the details below.
1.1. Motivation on Weyl’s law
For a bounded domain $\Omega\subset\mathbb{R}^{d}$ with Dirichlet boundary conditions, H. Weyl [42] found that the eigenvalue counting function $N(\lambda)$ for the Laplacian $\Delta$ below $\lambda$ satisfies the asymptotic formula:
$N(\lambda)=\frac{\omega_{d}{\rm vol}(\Omega)}{(2\pi)^{d}}\lambda^{\frac{d}{2}}+o(\lambda^{\frac{d}{2}}),\quad\mbox{as}\;\lambda\rightarrow\infty,$
where $\omega_{d}=\frac{\pi^{\frac{d}{2}}}{\Gamma(\frac{d}{2}+1)}$ is the volume of the unit ball in $\mathbb{R}^{d}$, and ${\rm vol}(\Omega)$ is the volume of $\Omega$. This result, now known as Weyl’s law, was later extended by Birman–Solomyak [1] to negative-order pseudo-differential operators in local settings of $\mathbb{R}^{d}$, and generalized to compact Riemannian manifolds by Chazarain [6] (see also [5, page 155]). In recent years, Weyl’s law has been actively investigated from both geometric [30, 15, 10, 7] and $C^{\ast}$-algebraic [27, 12, 36, 13] perspectives, with ongoing advances in results and associated techniques.
A powerful modern reformulation of Weyl’s law is expressed in terms of singular value asymptotics. Let $X$ be a $d$-dimensional compact smooth Riemannian manifold without boundary, and let $T$ be a pseudo-differential operator of order p-p on $X$ with $p>0$. In [14, Lemma 4.5], Grubb states (she referred to [1] without proof) that
(1.1) $\lim_{k\rightarrow\infty}k^{\frac{p}{d}}\mu(k,T)=\frac{1}{d(2\pi)^{d}}\|\sigma_{T}\|_{L_{\frac{d}{p}}(S^{*}X)}$
where $\sigma_{T}$ is the principal symbol of $T$ and $S^{\ast}X$ is the cosphere bundle of $X$. The formula (1.1) is related to Connes’s original noncommutative integration [2, 32, 31]. It is generalised to Theorem A on Euclidean space [12, 13] and a counterpart of Theorem A (with the case $p=d$) on quantum tori [36] including the classical tori.
1.2. Connection with noncommutative integration
Let $(X,G)$ be a compact smooth Riemannian manifold of dimension $d$, equipped with Laplace-Beltrami operator $\Delta_{G}$. In noncommutative geometry [2], A. Connes observed that Weyl’s law could be reinterpreted as follows: the integral of $f\in C^{\infty}(X)$ equals the Dixmier trace of operator $M_{f}(1+\Delta_{G})^{-\frac{d}{2}}$. In [2], the functional
${\rm Int}(T):A\mapsto{\rm Tr}_{\omega}\left(T(1+\Delta_{G})^{-\frac{d}{2}}\right),\quad T\in B(L_{2}(X,{\rm vol}_{G})),$
where ${\rm Tr}_{\omega}$ denote the Dixmier trace with respect to an extended limit $\omega$ on $\ell_{\infty}$, is proposed as a notion of noncommutative integration. Whenever $T(1+\Delta_{G})^{-\frac{d}{2}}$ belongs to $\mathcal{L}_{1,\infty}(L_{2}(X,{\rm vol}_{G}))$, this definition is independent of the choice of the extended limit $\omega$ and yields the same value (see [24] for more details). The connection between asymptotic formulas for singular values and noncommutative integral formulas was first studied by Rozenblum [32], and later explored in [38].
According to [22, Theorem 1.5] (see also [26, 24] for the case of classical tori), for any $S\in\Pi_{X}$, one has $T(1+\Delta_{G})^{-\frac{d}{2}}\in\mathcal{L}_{1,\infty}(L_{2}(X,{\rm vol}_{G}))$ and
${\rm Int}(S)=c_{d}\int_{T^{\ast}X}{\rm sym}_{X}(S)e^{-q_{G}}d\lambda.$
If $S\in\Pi_{X}$ is positive, Theorem 1.2 states that
${\rm Int}(S)=\lim_{t\rightarrow\infty}t\mu(t,S(1+\Delta_{G})^{-\frac{d}{2}}).$
Hence, Theorem 1.2 shows a direct and explicit link between noncommutative integral formula in [22] and singular value asymptotics for positive elements in $\Pi_{X}$.
1.3. Advantages and obstacles
In contrast to (1.1), the advantage of Theorem 1.2 is threefold: (a) it works on compact smooth Riemannian manifolds, including those with boundary; (b) it encompasses a strictly larger class of operators; (c) it is situated within the category of $C^{\ast}$-algebras (i.e. non-commutative topological spaces), not in a category of classical pseudodifferential operators which does not have a natural counterpart in non-commutative geometry. In addition, Theorem 1.2 covers a counterpart of Theorem A on classical tori [36] and provides Connes’ noncommutative integral for positive operators in $\Pi_{X}$.
Several significant obstacles arise to treat Theorem 1.2. First, since the Laplace–Beltrami operator locally equals to an elliptic differential operator, we need to establish spectral estimates for commutators involving elliptic differential operators on $\mathbb{R}^{d}$. Second, we must adapt Theorem A to a curved space setting with respect to an elliptic differential operator on $\mathbb{R}^{d}$. Third, transferring singular value asymptotics from local to global settings requires avoiding the Complex Powers Theorem for pseudo-differential operators on manifolds, as no satisfactory version is available.
Our proof diverges from the methods of Birman and Solomyak [1, 14] by employing a novel combination of functional calculus and double operator integrals. Two key innovations are the differentiability of operators inspiring by Connes–Moscovici (Subsection 5.2) and the new concept of “good atlas” on compact manifolds (Subsection 6.1). We demonstrate the density of differentiable elements in $\Pi_{X}$ (Lemma 5.9) and the existence of “good atlas” for any compact Riemannian manifold (Proposition 6.3), which serve as foundations to show Theorem 1.2. Collectively, these strategies go beyond the classical pseudo-differential calculus.
1.4. Outline of the Paper
Section 2 provides necessary notions and facts.
In Section 3, we establish spectral estimates (Theorems 3.1 and 3.2) for commutators of elliptic differential operators on $\mathbb{R}^{d}$ arising from the Laplace–Beltrami operator on manifolds.
In Section 4, we generalize Theorem A to Theorem 4.1 by replacing the Laplacian with an elliptic differential operator on $\mathbb{R}^{d}$ for the index $p=\frac{d}{2}$.
In Section 5, we give spectral characterization (Theorem 5.1) for powers of two nearly commutating operators.
In Section 6, we show Theorem 1.2 in a local chart by using the spectral estimates from Section 3 and the singular value asymptotics from Section 4.
Finally, in Section 7, we combine the spectral estimates from Sections 3 and 5 with the local version of Theorem 1.2 from Section 6 to end the proof.
Throughout, we use $c_{A}$ to denote some constant depending only on the symbol $A$, which may change from line to line without specification.
2. Preliminaries
In this section, we recall necessary notions for manifold and double operator integral, and provide some facts related to Cwikel estimate and asymptotics.
2.1. Compact manifold
The following materials are standard, see e.g. [4, 25]. Let $(X,G)$ be a $d$-dimensional compact smooth Riemannian manifold with Riemannian metric $G$ and an atlas $\{({\mathcal{U}}_{i},h_{i})\}_{i\in\mathbb{I}}$. Let
$\Omega_{i}=h_{i}({\mathcal{U}}_{i})\subset\mathbb{R}^{d},\quad\Omega_{i,j}=h_{i}({\mathcal{U}}_{i}\cap{\mathcal{U}}_{j})\subset\mathbb{R}^{d}.$
Denote $\Phi_{i,j}:\Omega_{i,j}\rightarrow\Omega_{j,i}$ the diffeomorphism given by the formula $\Phi_{i,j}=h_{j}\circ h_{i}^{-1}.$ Let $T^{\ast}X$ be the cotangent bundle of $X$ and let $\pi:T^{\ast}X\rightarrow X$ be the canonical projection. There exists an atlas $\{(\pi^{-1}({\mathcal{U}}_{i}),H_{i})\}_{i\in\mathbb{I}}$ on $T^{\ast}X$ such that
  1. (a)
    for $i\in\mathbb{I}$, function $H_{i}:\pi^{-1}({\mathcal{U}}_{i})\rightarrow\Omega_{i}\times\mathbb{R}^{d}$ is a homeomorphism;
  2. (b)
    for $i,j\in\mathbb{I}$ such that $\Omega_{i,j}\neq\emptyset$, we have (H_j∘H_i^-1)(ξ,s) =(Φ_i,j(ξ),(J_Φ_i,j^∗(ξ))^-1s), ξ∈Ω_i,j, s∈R^d, where $J_{\Phi_{i,j}}$ is the Jacobian matrix of mapping $\Phi_{i,j}$.
Since $T^{\ast}X$ is a smooth $d$-dimensional manifold, it follows that $T^{\ast}X$ has a canonical symplectic structure [16]. The corresponding Liouville measure $d\lambda$ on $T^{\ast}X$ is given by
$\int_{T^{\ast}X}(f\circ H_{i})d\lambda=\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}f(\xi,s)dm(\xi,s),\quad f\in C_{c}(\Omega_{i}\times\mathbb{R}^{d}),\;i\in\mathbb{I},$
where $dm$ is the product-Lebesgue measure on $\mathbb{R}^{d}\times\mathbb{R}^{d}$. Here $f\circ H_{i}$ denotes a function on $T^{\ast}X$ which equals $f\circ H_{i}$ on $\pi^{-1}({\mathcal{U}}_{i})$ and which vanishes outside $\pi^{-1}({\mathcal{U}}_{i})$. By the way, there is no canonical way to equip the cosphere bundle $S^{\ast}X$ of a smooth manifold $X$ with a measure.
Let $\Omega\subset\mathbb{R}^{d}$ be connected and open. Let $g:\Omega\rightarrow{\rm GL}^{+}(d,\mathbb{R})$ be a smooth mapping where ${\rm GL}^{+}(d,\mathbb{R})$ stands for the set of all positive elements in ${\rm GL}(d,\mathbb{R})$ (the space of all $d\times d$ invertible real matrices). The Laplace-Beltrami operator $\Delta_{g}:C_{c}^{\infty}(\Omega)\rightarrow C_{c}^{\infty}(\Omega)$ associated to mapping $g$ is defined by
$\Delta_{g}=M_{{\rm det}(g)^{-\frac{1}{2}}}\sum_{k,l=1}^{d}D_{k}M_{{\rm det}(g)^{\frac{1}{2}}\cdot(g^{-1})_{kl}}D_{l}$
where $D_{k}=\frac{1}{i}\partial_{k}$ and ${\rm det}(g)$ is the determinant of $g$.
For $i\in\mathbb{I}$, the components of the metric $G$ in the chart $({\mathcal{U}}_{i},h_{i})$ give rise to a smooth mapping $G_{i}:{\mathcal{U}}_{i}\rightarrow{\rm GL}^{+}(d,\mathbb{R})$. Set $g_{i}=G_{i}\circ h_{i}^{-1}:\Omega_{i}\rightarrow{\rm GL}^{+}(d,\mathbb{R})$. Define a partial isometry $W_{h_{i}}:L_{2}(X,{\rm vol}_{G})\to L_{2}(\mathbb{R}^{d},{\rm vol}_{g_{i}})$ by setting
$W_{h_{i}}\xi=\xi\circ h_{i}^{-1}.$
Clearly,
$W_{h_{i}}^{\ast}W_{h_{i}}=M_{\chi_{{\mathcal{U}}_{i}}},\quad W_{h_{i}}W_{h_{i}}^{\ast}=M_{\chi_{h_{i}({\mathcal{U}}_{i})}}.$
Since $X$ is compact, there is a good partition of unity $\{\phi_{n}\}_{n=1}^{N}\subset C_{c}^{\infty}(X)$ which is a finite partition of unity on $X$ and each $\phi_{n}$ is compactly supported in some chart. The Laplace-Beltrami operator $\Delta_{G}:C^{\infty}(X)\rightarrow C^{\infty}(X)$ is defined by the formula
$\Delta_{G}f=\sum_{n=1}^{N}\Delta_{g_{i_{n}}}\Big((f\phi_{n})\circ h_{i_{n}}^{-1}\Big)\circ h_{i_{n}},\quad f\in C^{\infty}(X).$
Here $i_{n}$ is chosen such that $\phi_{n}$ is compactly supported in ${\mathcal{U}}_{i_{n}}$. This definition does not rely on the choice of the good partition of unity $\{\phi_{n}\}_{n=1}^{N}$.
For our purpose, we need to introduce the canonical weight $e^{-q_{G}}$ on $T^{\ast}X$. Set
$q_{i}(\xi,s)=\langle g_{i}(\xi)^{-1}s,s\rangle,\quad\xi\in\Omega_{i},\;s\in\mathbb{R}^{d}.$
By [22, Fact 2.10], there exists a function $q_{G}$ defined on $T^{\ast}X$ such that
$q_{i}=q_{G}\circ H_{i}^{-1}\quad\mbox{on}\quad\Omega_{i}\times\mathbb{R}^{d}.$
The function $q_{G}$ is the square of the length function on $T^{\ast}X$ defined by the induced Riemannian metric on the cotangent bundle $T^{\ast}X$. More exactly, in local coordinates, the function $e^{-q_{G}}$ satisfies $e^{-q_{G}}\circ H_{i}^{-1}(\xi,s)=e^{-q_{i}(\xi,s)}$ on $\Omega_{i}\times\mathbb{R}^{d}$. We call $e^{-q_{G}}$ the canonical weight of $(X,G)$ on $T^{\ast}X$. Since $X$ is compact, it follows $e^{-q_{G}}\in L_{1}(T^{\ast}X,d\lambda)$.
In what follows, we always suppose that the manifold in considering is smooth.
2.2. Pseudo-differential operator
For $a\in C^{\infty}(\mathbb{R}^{d}\times\mathbb{R}^{d})$, define
${\rm Op}(a)f(x)=\int_{\mathbb{R}^{d}}e^{2\pi{\rm i}\langle x,\xi\rangle}a(x,\xi)\hat{f}(\xi)d\xi,\quad f\in L_{2}(\mathbb{R}^{d}),$
where $\hat{f}$ is the Fourier transform of $f$. By the Calderón-Vaillancourt theorem, ${\rm Op}(a)$ is a bounded on $L_{2}(\mathbb{R}^{d})$. For $m\in\mathbb{Z}$, let $\Psi^{m}(\mathbb{R}^{d})$ be the class of operators ${\rm Op}(a)$ with its symbol $a$ satisfying
$|D_{x}^{\alpha}D_{\xi}^{\beta}a(x,\xi)|\leq c_{\alpha,\beta}(1+|\xi|_{2})^{\frac{m-|\beta|_{1}}{2}},\quad\alpha,\beta\in\mathbb{Z}_{+}^{d}.$
Here $|\beta|_{1}=\sum_{k=1}^{d}|\beta_{k}|$ and $|\xi|_{2}=(\sum_{k=1}^{d}|\xi_{k}|^{2})^{\frac{1}{2}}$. The elements in $\Psi^{m}(\mathbb{R}^{d})$ are called pseudo-differential operators of order $m$. The key properties are that
$\Psi^{m_{1}}(\mathbb{R}^{d})\cdot\Psi^{m_{2}}(\mathbb{R}^{d})\subset\Psi^{m_{1}+m_{2}}(\mathbb{R}^{d}),\quad[\Psi^{m_{1}}(\mathbb{R}^{d}),\Psi^{m_{2}}(\mathbb{R}^{d})]\subset\Psi^{m_{1}+m_{2}-1}(\mathbb{R}^{d})$
for $m_{1},m_{2}\in\mathbb{Z}$.
Let $(X,G)$ be a compact Riemannian manifold. A linear operator $A:C^{\infty}(X)\rightarrow C^{\infty}(X)$ is called a pseudo-differential operator on $X$ of order $m$ if, for any chart $(\mathcal{U},h)$ and any $\psi,\phi\in C_{c}^{\infty}(\mathcal{U})$, the operator $M_{\psi}AM_{\phi}:C^{\infty}(\mathcal{U})\rightarrow C^{\infty}(\mathcal{U})$ is a pseudo-differential operator on $\mathbb{R}^{d}$ of order $m$. Denote $\Psi^{m}(X)$ the class of all pseudo-differential operators on $X$ of order $m$. Clearly, $\Delta_{G}\in\Psi^{2}(X)$. Similarly,
$\Psi^{m_{1}}(X)\cdot\Psi^{m_{2}}(X)\subset\Psi^{m_{1}+m_{2}}(X),\quad[\Psi^{m_{1}}(X),\Psi^{m_{2}}(X)]\subset\Psi^{m_{1}+m_{2}-1}(X)$
for $m_{1},m_{2}\in\mathbb{Z}$.
More information on pseudo-differential operators, we refer [22, 43] for $\mathbb{R}^{d}$ and [44, 34] for compact manifolds.
2.3. Sobolev space
The Sobolev space $W^{m,2}(\mathbb{R}^{d})$, $m\in\mathbb{N}$, consists of all distributions $f\in L_{2}(\mathbb{R}^{d})$ such that distribution $(1+\Delta)^{\frac{m}{2}}f$ belongs to $L_{2}(\mathbb{R}^{d})$.
A linear differential operator
$P=\sum_{k,l=1}^{d}M_{a_{kl}}D_{k}D_{l}+\sum_{k=1}^{d}M_{b_{k}}D_{k}+M_{f},\quad a_{kl},b_{k},f\in L_{1,loc}(\mathbb{R}^{d})$
is said to be elliptic if $\sum_{k,l=1}^{d}a_{kl}(x)\xi_{k}\xi_{l}\neq 0$ for all $\xi_{k}\neq 0$ ($1\leq k\leq d$). The Sobolev spaces $W^{2,2}(\mathbb{R}^{d})$ is the domain of the positive operator $\Delta$. Moreover, we crucially use the following fundamental result related to elliptic differential operator which is organised from [44].
Theorem 2.1.
Let $P:W^{2,2}(\mathbb{R}^{d})\to L_{2}(\mathbb{R}^{d})$ be a positive, self-adjoint, elliptic differential operator whose coefficients are constant outside some ball. For every $\alpha>0,$ we have $(1+P)^{-\frac{\alpha}{2}}:L_{2}(\mathbb{R}^{d})\to W^{\alpha,2}(\mathbb{R}^{d})$ is bounded.
Let $(X,G)$ be a compact Riemannian manifold and let $\Delta_{G}$ be the related Laplace-Beltrami operator. For any $m\in\mathbb{N}$, a distribution $f\in L_{2}(X,{\rm vol}_{G})$ belongs to the Sobolev space $W^{m,2}(X,{\rm vol}_{G})$ if and only if
$(1+\Delta_{G})^{\frac{m}{2}}f\in L_{2}(X,{\rm vol}_{G})$
where the operator is defined spectrally. This global, operator-theoretic characterization is equivalent to the local coordinate-based definition via a good partition of unity, a consequence of the manifold’s compactness. If $P:C^{\infty}(X)\rightarrow C^{\infty}(X)$ is a pseudo-differential operator of order $m$, then $P$ has a continuous extension $P:W^{k,2}(X,{\rm vol}_{G})\rightarrow W^{k-m,2}(X,{\rm vol}_{G})$.
2.4. Cwikel estimates
Given a Hilbert space $H$, denote $B(H)$ the algebra of all bounded linear operators on $H$. The ideal of all compact linear operators is denoted by $\mathcal{K}(H)$. Given $T\in\mathcal{K}(H)$, the singular value sequence $\mu(T)=\{\mu(n,T)\}_{n=0}^{\infty}$ is defined in term of the singular value function
$\mu(t,T)=\inf\{\|T-R\|_{B(H)}:{\rm Rank}(R)\leq t\},\quad t\geq 0.$
When necessary, we write $\mu_{B(H)}(t,T)$ to emphasize the ambient algebra.
For $0<p<\infty$, the Schatten ideal $\mathcal{L}_{p}(H)$ consists of all compact operators $T$ such that $\mu(T)\in\ell_{p}$. In this paper, we concern on the weak Schatten ideal $\mathcal{L}_{p,\infty}(H)$, which consists of all compact operators $T$ with $\mu(T)\in\ell_{p,\infty}$ and (quasi)-norm
$\|T\|_{\mathcal{L}_{p,\infty}(H)}:=\sup_{n\geq 0}(n+1)^{\frac{1}{p}}\mu(n,T).$
For convenience, $B(H)$ is also denoted by $\mathcal{L}_{\infty}(H)$. One has the following Hölder-type inequality
$\|TS\|_{\mathcal{L}_{r,\infty}(H)}\leq c_{p,q,r}\|T\|_{\mathcal{L}_{p,\infty}(H)}\|T\|_{\mathcal{L}_{q,\infty}(H)},\quad\frac{1}{r}=\frac{1}{p}+\frac{1}{q},\quad p,q,r\geq 1.$
For the purpose, denote by $(\mathcal{L}_{p,\infty})_{0}(H)$ the separable part of $\mathcal{L}_{p,\infty}(H)$, defined as the subset of $T\in\mathcal{L}_{p,\infty}(H)$ such that
$\lim_{n\rightarrow\infty}(n+1)^{\frac{1}{p}}\mu(n,T)=0.$
A fact we will use repeatedly is that $\mathcal{L}_{p,\infty}(H)\subset(\mathcal{L}_{p+\epsilon,\infty})_{0}(H)$ for any $\epsilon>0$. For further details on trace ideals and the singular value function, see e.g. [19, 35, 24].
Let us give some relevant Cwikel estimates here on $\mathbb{R}^{d}$ and compact manifold.
Lemma 2.2.
Let $\alpha,\gamma\geq 0$ be such that $\alpha+\gamma>0$. If $f\in C_{c}^{\infty}(\mathbb{R}^{d})$, then
$(1+\Delta)^{-\frac{\alpha}{2}}M_{f}(1+\Delta)^{-\frac{\gamma}{2}}\in\mathcal{L}_{\frac{d}{\alpha+\gamma},\infty}(L_{2}(\mathbb{R}^{d})).$
Proof.
Write $f=f_{1}f_{2}$ with $f_{1},f_{2}\in C_{c}^{\infty}(\mathbb{R}^{d})$. By Hölder’s inequality, it suffices to show
$(1+\Delta)^{-\frac{\alpha}{2}}M_{f_{1}}\in\mathcal{L}_{\frac{d}{\alpha},\infty}(L_{2}(\mathbb{R}^{d})),\quad M_{f_{2}}(1+\Delta)^{-\frac{\gamma}{2}}\in\mathcal{L}_{\frac{d}{\gamma},\infty}(L_{2}(\mathbb{R}^{d})).$
Here $\mathcal{L}_{\frac{d}{\alpha},\infty}(L_{2}(\mathbb{R}^{d}))$ ($\mathcal{L}_{\frac{d}{\gamma},\infty}(L_{2}(\mathbb{R}^{d}))$) is understood as $\mathcal{L}_{\infty}(L_{2}(\mathbb{R}^{d}))$ if $\alpha=0$ ($\gamma=0$). The inclusions follow from [13, Theorem 2.4]. ∎
Lemma 2.3.
Let $(X,G)$ be a $d$-dimensional compact Riemannian manifold. For well-defined function $\psi$ on $X$, we have
$\|M_{\psi}(1+\Delta_{G})^{-\frac{d}{2p}}\|_{\mathcal{L}_{p,\infty}(L_{2}(X,{\rm vol}_{G}))}\leq c_{p,X,G}\|\psi\|_{L_{p}(X,{\rm vol}_{G})},\quad p>2,$
$\|M_{\psi}(1+\Delta_{G})^{-\frac{d}{2p}}\|_{\mathcal{L}_{p,\infty}(L_{2}(X,{\rm vol}_{G}))}\leq c_{p,X,G}\|\psi\|_{L_{2}(X,{\rm vol}_{G})},\quad 0<p<2,$
$\|M_{|\psi|^{\frac{1}{2}}}(1+\Delta_{G})^{-\frac{d}{2p}}\|_{\mathcal{L}_{p,\infty}(L_{2}(X,{\rm vol}_{G}))}\leq c_{p,X,G}\|\psi\|^{\frac{1}{2}}_{L_{\Phi}(X,{\rm vol}_{G})},\quad p=2,$
where $L_{\Phi}(X,{\rm vol}_{G})$ is the Orlicz space associated with the Young function $\Phi(t)=t\log(e+t)$, $t>0$.
Proof.
The hardest case $p=2$ is established in [38, Theorem 1.3]. The proof in cases $p>2$ and $p<2$ is virtually identical. One simply needs to replace special Cwikel estimates established in [39] with (a) Cwikel estimates for $p>2$ (b) Birman-Solomyak version of Cwikel estimates for $p<2$ — see [23]. ∎
2.5. Birman-Solomyak lemmas
Fix $H$ a separable Hilbert space. The operators $\{A_{k}\}_{k\geq 0}\subset B(H)$ are called pairwise orthogonal if $A_{k}A_{l}=A_{k}^{\ast}A_{l}=A_{k}A_{l}^{\ast}=0$ whenever $k\neq l.$ Let us recall the following lemmas.
Lemma 2.4.
[12, Lemma 3.3] Let $\{A_{k}\}_{k=1}^{K}\subset\mathcal{L}_{p,\infty}(H)$ be a sequence of pairwise orthogonal operators such that, for all $1\leq k\leq N$, the limits
$\lim_{t\rightarrow\infty}t^{\frac{1}{p}}\mu(t,A_{k})=c_{k}\quad\mbox{exist}.$
Then
$\lim_{t\rightarrow\infty}t^{\frac{1}{p}}\mu(t,\sum_{k=1}^{K}A_{k})=(\sum_{k=1}^{K}c_{k}^{p})^{\frac{1}{p}}.$
An important feature for asymptotics of $(\mathcal{L}_{p,\infty})_{0}$ is the following.
Lemma 2.5.
Let $0<p<\infty.$ Let $A\in\mathcal{L}_{p,\infty}(H)$ be such that
$\lim_{t\to\infty}t^{\frac{1}{p}}\mu(t,A)=c.$
For every $B\in(\mathcal{L}_{p,\infty})_{0}(H),$ we have
$\lim_{t\to\infty}t^{\frac{1}{p}}\mu(t,A+B)=c.$
The elementary fact is ideed due to K. Fan, see e.g. [19, Chapter II, Theorem 4.1]. Its a generalisation due to Birman and Solomyak is the following one.
Lemma 2.6.
[33, 12] Let $0<p<\infty$. Let $\{A_{n}\}_{n\geq 1}\subset\mathcal{L}_{p,\infty}(H)$ be such that
  1. (a)
    $A_{n}\rightarrow A$ in $\mathcal{L}_{p,\infty}(H)$ as $n\rightarrow\infty$;
  2. (b)
    for each $n\geq 1$, the limit $\lim_{t\rightarrow\infty}t^{\frac{1}{p}}\mu(t,A_{n})=c_{n}$ exists.
Then the following limits exist and are equal,
$\lim_{t\rightarrow\infty}t^{\frac{1}{p}}\mu(t,A)=\lim_{n\rightarrow\infty}c_{n}.$
These three lemmas will play important roles in the proof of Theorem 1.2.
3. Compactness of commutator on Euclidean space
In this section, we prove Theorems 3.1 and 3.2, the spectral estimates related to elliptic differential operator on $\mathbb{R}^{d}$ arising from the Laplace–Beltrami operator on manifold.
Theorem 3.1.
Let $P:W^{2,2}(\mathbb{R}^{d})\to L_{2}(\mathbb{R}^{d})$ be a positive, self-adjoint, elliptic differential operator whose coefficients are constant outside some ball. Let $\phi\in C^{\infty}_{c}(\mathbb{R}^{d})$. If $\alpha,\beta,\gamma\in\mathbb{R}$ satisfy $\alpha+\beta+\gamma+1>0$, then
$(1+P)^{-\frac{\alpha}{2}}[M_{\phi},(1+P)^{-\frac{\beta}{2}}](1+P)^{-\frac{\gamma}{2}}\in\mathcal{L}_{\frac{d}{\alpha+\gamma+\beta+1},\infty}(L_{2}(\mathbb{R}^{d})).$
Theorem 3.2.
Let $P:W^{2,2}(\mathbb{R}^{d})\to L_{2}(\mathbb{R}^{d})$ be a positive, self-adjoint, elliptic differential operator whose coefficients are constant outside some ball. Let $0<s<2$ and $0\leq\alpha,\gamma<\infty$ with $\alpha+\gamma+s+1>0$. If $\phi\in C_{c}^{\infty}(\mathbb{R}^{d})$, then
$(1+P)^{-\frac{\alpha}{2}}[M_{\phi}R_{j}M_{\phi},(1+P)^{-\frac{s}{2}}](1+P)^{-\frac{\gamma}{2}}\in\mathcal{L}_{\frac{d}{\alpha+\gamma+s+1},\infty}(L_{2}(\mathbb{R}^{d}))$
where $R_{j}:=\partial_{j}\Delta^{-\frac{1}{2}}$ is the Riesz transform on $\mathbb{R}^{d}$ for $j=1,\cdots,d$.
3.1. Double operator integral
In this subsection, we demand the notion of double operator integrals as specifically used in [17, 29]. In what follows, $H$ is a separable Hilbert space. For self-adjoint operators $A_{1},A_{2}$ on $H$, let’s formally define the operator $T_{\psi}^{A_{1},A_{2}}$ with $\psi\in L_{\infty}(\mathbb{R}^{2})$ by setting
$T_{\psi}^{A_{1},A_{2}}(x)=\int_{\mathbb{R}^{2}}\psi(\lambda,\mu)dE^{1}(\lambda)xdE^{2}(\mu),\quad x\in B(H),$
where $E^{1},E^{2}$ are the spectral measures of $A_{1},A_{2}$ separably. We have below result.
Lemma 3.3.
For $0<\beta<m$, define
$\psi_{\beta,m}(\lambda,\mu)=\frac{\lambda^{\beta}-\mu^{\beta}}{\lambda^{m}-\mu^{m}}\lambda^{\frac{m-\beta}{2}}\mu^{\frac{m-\beta}{2}},\quad\lambda,\mu>0.$
For any positive operator $B\in B(H)$, we have
$\|T_{\psi_{\beta,m}}^{B,B}\|_{B(H)\rightarrow B(H)}<\infty.$
Proof.
Let $\frac{\lambda}{\mu}=e^{t}$ with $t\in\mathbb{R}$. We have
$\psi_{\beta,m}(\lambda,\mu)=\frac{(\frac{\lambda}{\mu})^{\beta}-1}{(\frac{\lambda}{\mu})^{m}-1}(\frac{\lambda}{\mu})^{\frac{m-\beta}{2}}=\frac{\sinh(\frac{\beta}{2}t)}{\sinh(\frac{m}{2}t)}$
where $t\mapsto\sinh(t)$ is the hyperbolic sine function. As $\beta<m$, the function $t\mapsto\frac{\sinh(\frac{\beta}{2}t)}{\sinh(\frac{m}{2}t)}$ belongs to the Schwartz class of $\mathbb{R}$. Let $g_{\beta,m}$ be its Fourier transform. By Fourier inversion,
$\psi_{\beta,m}(\lambda,\mu)=\int_{\mathbb{R}}g_{\beta,m}(s)e^{2\pi{\rm{i}}s\log_{e}(\frac{\lambda}{\mu})}ds=\int_{\mathbb{R}}g_{\beta,m}(s)\lambda^{2\pi{\rm{i}}s}\mu^{-2\pi{\rm{i}}s}ds.$
Therefore, for $A\in B(H)$,
$T_{\psi_{\beta,m}}^{B,B}(A)=\int_{\mathbb{R}}g_{\beta,m}(s)B^{2\pi{\rm{i}}s}AB^{-2\pi{\rm{i}}s}ds.$
Since $g_{\beta,m}$ is integrable in $L_{1}(\mathbb{R})$, the boundedness of $T_{\psi_{\beta,m}}^{B,B}$ follows. ∎
The following lemma is a direct result from functional calculus. Its a special case $m=2$ is established in [28, lemma 5.4] so we omit the calculations.
Lemma 3.4.
Let $B,A\in B(H)$ and assume that there is a constant $c>0$ such that $c\leq B\leq c^{-1}$. Let $\alpha,\gamma\geq 0$, and $0<\beta<m$. Define $\psi_{s,m}$ as in Lemma 3.3. We have
$B^{-\alpha}[A,B^{\beta}]B^{-\gamma}=T_{\psi_{\beta,m}}^{B,B}(B^{\frac{\beta-m}{2}-\alpha}[A,B^{m}]B^{\frac{\beta-m}{2}-\gamma}).$
Corollary 3.5.
Let $A,B$ be positive operators on Hilbert space $H$ such that $A:{\rm dom}(B^{m})\rightarrow{\rm dom}(B^{m})$, $m>0$. Assume that $A$ is bounded on $H$ and ${\rm ker}(B)=\{0\}$. For $0<\beta<m$ and $\alpha,\gamma\in\mathbb{R}$, we have
$\|B^{-\alpha}[A,B^{\beta}]B^{-\gamma}\|_{B(H)}\leq c_{\beta,m}\|B^{\frac{\beta-m}{2}-\alpha}[A,B^{m}]B^{\frac{\beta-m}{2}-\gamma}\|_{B(H)}$
with constant $c_{\beta,m}$ depending on $\beta,m$.
Proof.
For each $n\geq 1$, consider projection $P_{n}=\chi_{(\frac{1}{n},n)}(B)$ and Hilbert space $H_{n}=P_{n}(H)$. Let $B_{n}=B|_{H_{n}}$ and $A_{n}=P_{n}AP_{n}|_{H_{n}}$. Clearly, $A_{n}\in B(H_{n})$ and $B_{n}\in B(H_{n})$ satisfies the conditions of Lemma 3.4. Hence,
$B_{n}^{-\alpha}[A_{n},B_{n}^{\beta}]B_{n}^{-\gamma}=T_{\psi_{\beta,m}}^{B_{n},B_{n}}(B_{n}^{\frac{\beta-m}{2}-\alpha}[A_{n},B_{n}^{m}]B_{n}^{\frac{\beta-m}{2}-\gamma}).$
Using Lemma 3.3,
$\|B_{n}^{-\alpha}[A_{n},B_{n}^{\beta}]B_{n}^{-\gamma}\|_{B(H)}\leq c_{\beta,m}\|B_{n}^{\frac{\beta-m}{2}-\alpha}[A_{n},B_{n}^{m}]B_{n}^{\frac{\beta-m}{2}-\gamma}\|_{B(H)}.$
Observe that
$B_{n}^{-\alpha}[A_{n},B_{n}^{\beta}]B_{n}^{-\gamma}=P_{n}\cdot B^{-\alpha}[A,B^{\beta}]B^{-\gamma}\cdot P_{n}$
and
$B_{n}^{\frac{\beta-m}{2}-\alpha}[A_{n},B_{n}^{m}]B_{n}^{\frac{\beta-m}{2}-\gamma}=P_{n}\cdot B^{\frac{\beta-m}{2}-\alpha}[A,B^{m}]B^{\frac{\beta-m}{2}-\gamma}\cdot P_{n}.$
Therefore,
$\|P_{n}\cdot B^{-\alpha}[A,B^{\beta}]B^{-\gamma}\cdot P_{n}\|_{B(H)}\leq c_{\beta,m}\|B^{\frac{\beta-m}{2}-\alpha}[A,B^{m}]B^{\frac{\beta-m}{2}-\gamma}\|_{B(H)}$
for every $n\geq 1$. The assertion follows from the Fatou property in $B(H)$. ∎
3.2. Proof of Theorem 3.1
Before proving Theorem 3.1, we establish the following necessary lemmas.
Lemma 3.6.
Let $P$ and $\phi$ be given as in Theorem 3.1. For every $n\in\mathbb{N},$ we have
$(1+P)^{-\frac{2n-1}{2}}M_{\phi}D_{k}(1+P)^{n-1},(1+P)^{n}M_{\phi}D_{k}(1+P)^{-\frac{2n+1}{2}}\in B(L_{2}(\mathbb{R}^{d})).$
This should be understood as follows: operators are well defined on $\bigcap_{k\geq 0}W^{2k,2}(\mathbb{R}^{d})$ and extend to bounded operators on $L_{2}(\mathbb{R}^{d}).$
Proof.
By the functional calculus,
$(1+P)^{\alpha}:\bigcap_{k\geq 0}{\rm dom}((1+P)^{k})\to\bigcap_{k\geq 0}{\rm dom}((1+P)^{k}),\quad z\in\mathbb{C}.$
Hence, by Theorem 2.1,
$(1+P)^{\alpha}:\bigcap_{k\geq 0}W^{2k,2}(\mathbb{R}^{d})\to\bigcap_{k\geq 0}W^{2k,2}(\mathbb{R}^{d}),\quad z\in\mathbb{C}.$
Clearly, also
$M_{\phi}:\bigcap_{k\geq 0}W^{2k,2}(\mathbb{R}^{d})\to\bigcap_{k\geq 0}W^{2k,2}(\mathbb{R}^{d}).$
Hence, the operators in question are well defined on $\cap_{k\geq 0}W^{2k,2}(\mathbb{R}^{d}).$
We write the second operator as
$\Big((1+P)^{n}M_{\phi}D_{k}(1+P)^{-\frac{2n+1}{2}}:L_{2}(\mathbb{R}^{d})\to L_{2}(\mathbb{R}^{d})\Big)$
$\Big((1+P)^{n}M_{\phi}D_{k}:W^{2n+1,2}(\mathbb{R}^{d})\to L_{2}(\mathbb{R}^{d})\Big)\circ$
$\circ\Big((1+P)^{-\frac{2n+1}{2}}:L_{2}(\mathbb{R}^{d})\to W^{2n+1,2}(\mathbb{R}^{d})\Big).$
The first factor on the right hand side is obviously bounded. The second factor on the right hand side is bounded by Theorem 2.1. Hence, our second operator is bounded.
For every $\xi,\eta\in\cap_{k\geq 0}W^{2k,2}(\mathbb{R}^{d}),$ we have
$\langle(1+P)^{-\frac{2n-1}{2}}M_{\phi}D_{k}(1+P)^{n-1}\xi,\eta\rangle=\langle(1+P)^{n-1}D_{k}M_{\bar{\phi}}(1+P)^{-\frac{2n-1}{2}}\xi,\eta\rangle.$
That the operator
$(1+P)^{n-1}D_{k}M_{\bar{\phi}}(1+P)^{-\frac{2n-1}{2}}$
is bounded can be seen as in the preceding paragraph. Thus, our first operator is also bounded. ∎
Lemma 3.7.
Let $P$ and $\phi$ be given as in Theorem 3.1. For every $z\in\mathbb{C},$ we have
$(1+P)^{-\frac{z}{2}}M_{\phi}D_{k}(1+P)^{\frac{z-1}{2}},(1+P)^{-\frac{z}{2}}M_{\phi}(1+P)^{\frac{z-1}{2}}\in B(L_{2}(\mathbb{R}^{d})).$
This should be understood as follows: operators are well defined on $\bigcap_{k\geq 0}W^{2k,2}(\mathbb{R}^{d})$ and extend to bounded operators on $L_{2}(\mathbb{R}^{d}).$
Proof.
By the functional calculus,
$(1+P)^{z}:\bigcap_{k\geq 0}{\rm dom}((1+P)^{k})\to\bigcap_{k\geq 0}{\rm dom}((1+P)^{k}),\quad z\in\mathbb{C}.$
Hence, by Theorem 2.1,
$(1+P)^{z}:\bigcap_{k\geq 0}W^{2k,2}(\mathbb{R}^{d})\to\bigcap_{k\geq 0}W^{2k,2}(\mathbb{R}^{d}),\quad z\in\mathbb{C}.$
Clearly, also
$M_{\phi}:\bigcap_{k\geq 0}W^{2k,2}(\mathbb{R}^{d})\to\bigcap_{k\geq 0}W^{2k,2}(\mathbb{R}^{d}).$
Hence, the operator in question is well defined on $\cap_{k\geq 0}W^{2k,2}(\mathbb{R}^{d}).$
For $\xi,\eta\in\cap_{k\geq 0}W^{2k,2}(\mathbb{R}^{d}),$ define entire function
$F_{\xi,\eta}:z\to\langle(1+P)^{-\frac{z}{2}}M_{\phi}D_{k}(1+P)^{\frac{z-1}{2}}\xi,\eta\rangle,\quad z\in\mathbb{C}.$
Fix $n\in\mathbb{N}$ and consider the function $F_{\xi,\eta}$ on the strip $\{-2n\leq\Re(z)\leq 2n-1\}.$ We have
$F_{\xi,\eta}(z)=\langle M_{\phi}D_{k}(1+P)^{\frac{z-1}{2}}\xi,(1+P)^{-\frac{\bar{z}}{2}}\eta\rangle.$
Hence,
$|F_{\xi,\eta}(z)|$
$\leq\Big\|M_{\phi}D_{k}(1+P)^{\frac{z-1}{2}}\xi\Big\|_{L_{2}(\mathbb{R}^{d})}\Big\|(1+P)^{-\frac{\bar{z}}{2}}\eta\Big\|_{L_{2}(\mathbb{R}^{d})}$
$\leq\Big\|M_{\phi}D_{k}\Big\|_{W^{1,2}(\mathbb{R}^{d})\to L_{2}(\mathbb{R}^{d})}\Big\|(1+P)^{\frac{z-1}{2}}\xi\Big\|_{W^{1,2}(\mathbb{R}^{d})}\Big\|(1+P)^{-\frac{\bar{z}}{2}}\eta\Big\|_{L_{2}(\mathbb{R}^{d})}$
[…]
$\leq c_{P}\Big\|M_{\phi}D_{k}\Big\|_{W^{1,2}(\mathbb{R}^{d})\to L_{2}(\mathbb{R}^{d})}\|\xi\|_{{\rm dom}((P+1)^{\frac{2n-1}{2}})}\|\eta\|_{{\rm dom}(P^{n})}$
[…]
Thus, $F_{\xi,\eta}$ is bounded on the strip $\{-2n\leq\Re(z)\leq 2n-1\}.$
Clearly,
$F_{\xi,\eta}(2n-1+it)=\langle(1+P)^{-\frac{2n-1}{2}}M_{\phi}D_{k}(1+P)^{n-1}\big((1+P)^{\frac{it}{2}}\xi\big),(1+P)^{\frac{it}{2}}\eta\rangle.$
We, therefore, have
$|F_{\xi,\eta}(2n-1+it)|$
$\leq\Big\|(1+P)^{-\frac{2n-1}{2}}M_{\phi}D_{k}(1+P)^{n-1}\big((1+P)^{\frac{it}{2}}\xi\big)\Big\|_{L_{2}(\mathbb{R}^{d})}\Big\|(1+P)^{\frac{it}{2}}\eta\Big\|_{L_{2}(\mathbb{R}^{d})}$
$\leq\Big\|(1+P)^{-\frac{2n-1}{2}}M_{\phi}D_{k}(1+P)^{n-1}\Big\|_{L_{2}(\mathbb{R}^{d})\to L_{2}(\mathbb{R}^{d})}\times$
$\qquad\times\Big\|(1+P)^{\frac{it}{2}}\xi\Big\|_{L_{2}(\mathbb{R}^{d})}\Big\|(1+P)^{\frac{it}{2}}\eta\Big\|_{L_{2}(\mathbb{R}^{d})}$
$\leq\Big\|(1+P)^{-\frac{2n-1}{2}}M_{\phi}D_{k}(1+P)^{n-1}\Big\|_{L_{2}(\mathbb{R}^{d})\to L_{2}(\mathbb{R}^{d})}\|\xi\|_{L_{2}(\mathbb{R}^{d})}\|\eta\|_{L_{2}(\mathbb{R}^{d})}.$
Here, the finiteness of the right hand side is guaranteed by Lemma 3.6.
Similarly,
$|F_{\xi,\eta}(-2n+it)|$
$\leq\Big\|(1+P)^{n}M_{\phi}D_{k}(1+P)^{-\frac{2n+1}{2}}\Big\|_{L_{2}(\mathbb{R}^{d})\to L_{2}(\mathbb{R}^{d})}\|\xi\|_{L_{2}(\mathbb{R}^{d})}\|\eta\|_{L_{2}(\mathbb{R}^{d})}.$
By the Maximum Modulus Principle,
$|F_{\xi,\eta}(z)|$
$\leq\Big(\Big\|(1+P)^{-\frac{2n-1}{2}}M_{\phi}D_{k}(1+P)^{n-1}\Big\|_{L_{2}(\mathbb{R}^{d})\to L_{2}(\mathbb{R}^{d})}+$
$+\Big\|(1+P)^{n}M_{\phi}D_{k}(1+P)^{-\frac{2n+1}{2}}\Big\|_{L_{2}(\mathbb{R}^{d})\to L_{2}(\mathbb{R}^{d})}\Big)\times\|\xi\|_{L_{2}(\mathbb{R}^{d})}\|\eta\|_{L_{2}(\mathbb{R}^{d})}$
for every $z$ in the strip $\{-2n\leq\Re(z)\leq 2n-1\}.$ Hence,
$\Big\|(1+P)^{-\frac{z}{2}}M_{\phi}D_{k}(1+P)^{\frac{z-1}{2}}\Big\|_{L_{2}(\mathbb{R}^{d})\to L_{2}(\mathbb{R}^{d})}$
$\leq\Big\|(1+P)^{-\frac{2n-1}{2}}M_{\phi}D_{k}(1+P)^{n-1}\Big\|_{L_{2}(\mathbb{R}^{d})\to L_{2}(\mathbb{R}^{d})}+$
$+\Big\|(1+P)^{n}M_{\phi}D_{k}(1+P)^{-\frac{2n+1}{2}}\Big\|_{L_{2}(\mathbb{R}^{d})\to L_{2}(\mathbb{R}^{d})}$
for every $z$ in the strip $\{-2n\leq\Re(z)\leq 2n-1\}.$ This completes the proof. ∎
Lemma 3.8.
Let $P$ and $\phi$ be given as in Theorem 3.1. If $\alpha,\gamma\in\mathbb{R}$ satisfy $\alpha+\gamma=1$, then
$(1+P)^{-\frac{\alpha}{2}}[M_{\phi},P](1+P)^{-\frac{\gamma}{2}}\in B(L_{2}(\mathbb{R}^{d})).$
Proof.
Write
$P=\sum_{k,l=1}^{d}M_{a_{kl}}D_{k}D_{l}+\sum_{k=1}^{d}M_{b_{k}}D_{k}+M_{f},$
where functions $\{a_{kl}\}_{k,l=1}^{d},$ $\{b_{k}\}_{k=1}^{d}$ and $f$ are smooth and constant outside some ball. We have
$[M_{\phi},P]$ $=\sum_{k,l=1}^{d}M_{a_{kl}}[M_{\phi},D_{k}D_{l}]+\sum_{k=1}^{d}M_{b_{k}}[M_{\phi},D_{k}]$
$=-\sum_{k,l=1}^{d}M_{a_{kl}}\cdot M_{D_{k}\phi}D_{l}-\sum_{k,l=1}^{d}M_{a_{kl}}\cdot D_{k}M_{D_{l}\phi}-\sum_{k=1}^{d}M_{b_{k}}\cdot M_{D_{k}\phi}$
$=-\sum_{k,l=1}^{d}M_{a_{kl}}\cdot M_{D_{k}\phi}D_{l}-\sum_{k,l=1}^{d}M_{a_{kl}}\cdot M_{D_{l}\phi}D_{k}-\sum_{k,l=1}^{d}M_{a_{kl}}\cdot M_{D_{k}D_{l}\phi}+$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad-\sum_{k=1}^{d}M_{b_{k}}\cdot M_{D_{k}\phi}.$
The assertion follows now from Lemma 3.7. ∎
Lemma 3.9.
Let $\alpha,\beta\in\mathbb{R}.$ Set
$T_{\alpha,\beta}=(1+P)^{-\frac{\alpha}{2}}[M_{\phi},(1+P)^{\frac{\beta}{2}}](1+P)^{\frac{\alpha-\beta+1}{2}}.$
We have
$T_{\alpha,\beta+1}=T_{\alpha,\beta}+T_{\alpha-\beta,1}.$
Proof.
By the Leibniz rule,
$T_{\alpha,\beta+1}=(1+P)^{-\frac{\alpha}{2}}[M_{\phi},(1+P)^{\frac{\beta}{2}}\cdot(1+P)^{\frac{1}{2}}](1+P)^{\frac{\alpha-\beta}{2}}$
$=(1+P)^{-\frac{\alpha}{2}}[M_{\phi},(1+P)^{\frac{\beta}{2}}](1+P)^{\frac{\alpha-\beta+1}{2}}+(1+P)^{\frac{\beta-\alpha}{2}}[M_{\phi},(1+P)^{\frac{1}{2}}](1+P)^{\frac{\alpha-\beta}{2}}$
$=T_{\alpha,\beta}+T_{\alpha-\beta,1}.$
The assertion follows. ∎
Lemma 3.10.
Let $P$ and $\phi$ be given as in Theorem 3.1. If $\alpha,\beta,\gamma\in\mathbb{R}$ satisfy $\alpha+\gamma+1=\beta$, then
$(1+P)^{-\frac{\alpha}{2}}[M_{\phi},(1+P)^{\frac{\beta}{2}}](1+P)^{-\frac{\gamma}{2}}\in B(L_{2}(\mathbb{R}^{d})).$
Proof.
For $\beta=0,$ there is nothing to prove.
Case 1: Let $0\leq\beta<2.$ For $\beta=0,$ there is nothing to prove and, so, we assume that $0<\beta<2.$ Set $A=M_{\phi}$ and $B=(1+P)^{\frac{1}{2}}.$ By Corollary 3.5 and Lemma 3.8,
$\|(1+P)^{-\frac{\alpha}{2}}[M_{\phi},(1+P)^{\frac{\beta}{2}}](1+P)^{-\frac{\gamma}{2}}\|_{B(L_{2}(\mathbb{R}^{d}))}$
$\leq c_{\beta}\|(1+P)^{\frac{\beta-2}{4}-\frac{\alpha}{2}}[M_{\phi},P](1+P)^{\frac{\beta-2}{4}-\frac{\gamma}{2}}\|_{B(L_{2}(\mathbb{R}^{d}))}<\infty.$
This proves the assertion for $0<\beta<2.$
Case 2: Consider now the general case. By Lemma 3.9, we have
$T_{\alpha,\beta+1}=T_{\alpha,\beta}+T_{\alpha-\beta,1},\quad T_{\alpha,\beta-1}=T_{\alpha,\beta}-T_{\alpha-\beta+1,1}.$
By induction, we can express $T_{\alpha,\beta}$ as linear combination of $T_{\delta,\gamma}$’s with $\gamma\in[0,1].$ The assertion in general case follows now from that in Case 1. ∎
Lemma 3.11.
Let $P$ and $\phi$ be given as in Theorem 3.1. If $\alpha,\gamma\geq 0$ with $\alpha+\gamma>0$, then
$(1+P)^{-\frac{\alpha}{2}}M_{\phi}(1+P)^{-\frac{\gamma}{2}}\in\mathcal{L}_{\frac{d}{\alpha+\gamma},\infty}(L_{2}(\mathbb{R}^{d})).$
Proof.
Write
$(1+P)^{-\frac{\alpha}{2}}M_{\phi}(1+P)^{-\frac{\gamma}{2}}=(1+P)^{-\frac{\alpha}{2}}(1+\Delta)^{\frac{\alpha}{2}}\cdot(1-\Delta)^{-\frac{\alpha}{2}}M_{\phi}(1+\Delta)^{-\frac{\gamma}{2}}\cdot$
$\cdot(1+\Delta)^{\frac{\gamma}{2}}(1+P)^{-\frac{\gamma}{2}}.$
The first and the third factors on the right hand side are bounded by Theorem 2.1. The second factor on the right hand side belongs to $\mathcal{L}_{\frac{d}{\alpha+\gamma},\infty}(L_{2}(\mathbb{R}^{d}))$ by Lemma 2.2. ∎
It is time to prove Theorem 3.1.
Proof of Theorem 3.1.
Write $\phi=\phi_{1}\phi_{2}$ with $\phi_{1},\phi_{2}\in C_{c}^{\infty}(\mathbb{R}^{d})$. We have
$(1+P)^{-\frac{\alpha}{2}}[M_{\phi},(1+P)^{-\frac{\beta}{2}}](1+P)^{-\frac{\gamma}{2}}$
$=(1+P)^{-\frac{\alpha}{2}}M_{\phi_{1}}(1+P)^{-\frac{\beta+\gamma+1}{2}}\cdot(1+P)^{\frac{\beta+\gamma+1}{2}}[M_{\phi_{2}},(1+P)^{-\frac{\beta}{2}}](1+P)^{-\frac{\gamma}{2}}+$
$+(1+P)^{-\frac{\alpha}{2}}[M_{\phi_{1}},(1+P)^{-\frac{\beta}{2}}](1+P)^{\frac{\beta+\alpha+1}{2}}\cdot(1+P)^{-\frac{\beta+\alpha+1}{2}}M_{\phi_{2}}(1+P)^{-\frac{\gamma}{2}}.$
In the first summand on the right hand side, the first factor belongs to the ideal $\mathcal{L}_{\frac{1}{\alpha+\beta+\gamma+1},\infty}(L_{2}(\mathbb{R}^{d}))$ by Lemma 3.11, while the second factor is bounded by Lemma 3.10. In the second summand on the right hand side, the first factor is bounded by Lemma 3.10, while the second factor belongs to $\mathcal{L}_{\frac{1}{\alpha+\beta+\gamma+1},\infty}(L_{2}(\mathbb{R}^{d}))$ by Lemma 3.11. This completes the proof. ∎
3.3. Proof of Theorem 3.2
In this subsection, we prove Theorem 3.2.
Lemma 3.12.
Let $P$ and $\phi$ be given in Theorem 3.2. Given a smooth function $f$ on $\mathbb{R}^{d}$ and $0\leq s\leq 2$, we have
$(1+P)^{-\frac{s}{4}}[M_{f}D_{l},M_{\phi}R_{j}M_{\phi}](1+P)^{\frac{s}{4}-\frac{1}{2}}\in B(L_{2}(\mathbb{R}^{d})).$
Proof.
Write $\phi=\phi\theta$ with $\theta\in C_{c}^{\infty}(\mathbb{R}^{d})$. We have $f\theta\in C_{c}^{\infty}(\mathbb{R}^{d})$. Write
$[M_{f}D_{l},M_{\phi}R_{j}M_{\phi}]=[M_{f\theta}D_{l},M_{\phi}R_{j}M_{\phi}].$
Clearly, $M_{f\theta}D_{l}\in\Psi^{1}(\mathbb{R}^{d})$ and $M_{\phi}R_{j}M_{\phi}\in\Psi^{0}(\mathbb{R}^{d})$. We have $M_{f\theta}D_{l},M_{\phi}R_{j}M_{\phi}]\in\Psi^{0}(\mathbb{R}^{d})$ and therefor it is bounded on $L_{2}(\mathbb{R}^{d})$ by [43, Theorem 2,page 286]. Since $(1+P)^{-\frac{s}{4}}$ and $(1+P)^{\frac{s}{4}-\frac{1}{2}}$ are bounded for $0\leq s\leq 2$, the assertion follows. ∎
Lemma 3.13.
Let $P$ and $\phi$ be given in Theorem 3.2. Given a smooth function $f$ on $\mathbb{R}^{d}$ and $0\leq s\leq 2$, we have
$(1+P)^{-\frac{s}{4}}[M_{f}D_{i}D_{k},M_{\phi}R_{j}M_{\phi}](1+P)^{\frac{s}{4}-\frac{1}{2}}\in B(L_{2}(\mathbb{R}^{d})).$
Proof.
Since $\phi\in C_{c}^{\infty}(\mathbb{R}^{d})$, there is a function $\theta\in C_{c}^{\infty}(\mathbb{R}^{d})$ such that $\phi=\phi\theta$ and
$M_{f}D_{i}D_{k},M_{\phi}R_{j}M_{\phi}]=[M_{f\theta}D_{i}D_{k},M_{\phi}R_{j}M_{\phi}].$
We have $M_{f\theta}D_{i}D_{k}\in\Psi^{2}(\mathbb{R}^{d})$ and therefore $M_{f\theta}D_{i}D_{k},M_{\phi}R_{j}M_{\phi}]\in\Psi^{1}(\mathbb{R}^{d})$. Observe that
$M_{f\theta}D_{i}D_{k},M_{\phi}R_{j}M_{\phi}]=M_{\theta}[M_{f\theta}D_{i}D_{k},M_{\phi}R_{j}M_{\phi}]M_{\theta}.$
Thus,
$(1+\Delta)^{-\frac{s}{4}}[M_{f}D_{i}D_{k},M_{\phi}R_{j}M_{\phi}](1+\Delta)^{\frac{s}{4}-\frac{1}{2}}$
$=(1+\Delta)^{-\frac{s}{4}}M_{\theta}\cdot[M_{f\theta}D_{i}D_{k},M_{\phi}R_{j}M_{\phi}]\cdot M_{\theta}(1+\Delta)^{\frac{s}{4}-\frac{1}{2}}$
$\in\Psi^{-\frac{s}{2}}(\mathbb{R}^{d})\cdot\Psi^{1}(\mathbb{R}^{d})\cdot\Psi^{\frac{s}{2}-1}(\mathbb{R}^{d})$
$\subset\Psi^{0}(\mathbb{R}^{d}).$
Combining this fact with the boundedness of operators $(1+P)^{-\frac{s}{4}}(1+\Delta)^{\frac{s}{4}}$ and $(1+\Delta)^{\frac{s}{4}-\frac{1}{2}}(1+P)^{\frac{s}{4}-\frac{1}{2}}$, [43, Theorem 2,page 286] gives the assertion. ∎
Lemma 3.14.
Let $P$ and $\phi$ be given in Theorem 3.2. If $s\in(0,2)$, then
$(1+P)^{\frac{1-s}{2}}[(1+P)^{\frac{s}{2}},M_{\phi}R_{j}M_{\phi}]\in B(L_{2}(\mathbb{R}^{d})).$
Proof.
Using Corollary 3.5, there is a constant $c_{s}$ relying on $s$ such that
$\|(1+P)^{\frac{1-s}{2}}[(1+P)^{\frac{s}{2}},M_{\phi}R_{j}M_{\phi}]\|_{\infty}\leq c_{s}\|(1+P)^{-\frac{s}{4}}[P,M_{\phi}R_{j}M_{\phi}](1+P)^{\frac{s}{4}-\frac{1}{2}}\|_{\infty}.$
It suffices to prove that the factor on the right hand side is finite. Write
$P=M_{f_{0}}+\sum_{l=1}^{d}M_{f_{l}}D_{l}+\sum_{i,k=1}^{d}M_{f_{ik}}D_{i}D_{k},$
where functions $\{f_{ik}\}_{i,k=1}^{d},$ $\{f_{l}\}_{l=1}^{d}$ and $f_{0}$ are smooth and constant outside some ball on $\mathbb{R}^{d}$. Thus,
$(1+P)^{-\frac{s}{4}}[P,M_{\phi}R_{j}M_{\phi}](1+P)^{\frac{s}{4}-\frac{1}{2}}$
$=(1+P)^{-\frac{s}{4}}[M_{f_{0}},M_{\phi}R_{j}M_{\phi}](1+P)^{\frac{s}{4}-\frac{1}{2}}+$
$+\sum_{l=1}^{d}(1+P)^{-\frac{s}{4}}[M_{f_{l}}D_{l},M_{\phi}R_{j}M_{\phi}](1+P)^{\frac{s}{4}-\frac{1}{2}}+$
$+\sum_{i,k=1}^{d}(1+P)^{-\frac{s}{4}}[M_{f_{ik}}D_{i}D_{k},M_{\phi}R_{j}M_{\phi}](1+P)^{\frac{s}{4}-\frac{1}{2}}.$
By Lemma 3.12, the second factor on the right hand side is bounded. By Lemma 3.13, the third factor on the right hand side is bounded. Since $\phi\in C_{c}^{\infty}(\mathbb{R}^{d})$, write $\phi=\phi\psi$ with $\psi\in C_{c}^{\infty}(\mathbb{R}^{d})$. By $f_{0}\psi\in C_{c}^{\infty}(\mathbb{R}^{d})$, we infer from [9] that
$M_{f_{0}},M_{\phi}R_{j}M_{\phi}]=M_{\phi}[M_{f_{0}\psi},R_{j}]M_{\phi}\in B(L_{2}(\mathbb{R}^{d}))$
and therefore, for $s\in(0,2)$,
$(1+P)^{-\frac{s}{4}}[M_{f_{0}},M_{\phi}R_{j}M_{\phi}](1+P)^{\frac{s}{4}-\frac{1}{2}}\in B(L_{2}(\mathbb{R}^{d})).$
This is the estimate for the first factor and the proof is complete. ∎
Let us start the proof of Theorem 3.2.
Proof of Theorem 3.2.
Take $\psi\in C_{c}^{\infty}(\mathbb{R}^{d})$ such that $\phi=\phi\psi$ and therefore $M_{\phi}R_{j}M_{\phi}=M_{\psi}\cdot M_{\phi}R_{j}M_{\phi}\cdot M_{\psi}.$ We have
$(1+P)^{-\frac{\alpha}{2}}[M_{\phi}R_{j}M_{\phi},(1+P)^{-\frac{s}{2}}](1+P)^{-\frac{\gamma}{2}}$
$=(1+P)^{-\frac{\alpha}{2}}[M_{\psi},(1+P)^{-\frac{s}{2}}]\cdot M_{\phi}R_{j}\cdot M_{\phi}(1+P)^{-\frac{\gamma}{2}}+$
$+(1+P)^{-\frac{\alpha}{2}}M_{\psi}[M_{\phi}R_{j}M_{\phi},(1+P)^{-\frac{s}{2}}]M_{\psi}(1+P)^{-\frac{\gamma}{2}}+$
$+(1+P)^{-\frac{\alpha}{2}}M_{\phi}\cdot R_{j}M_{\phi}\cdot[M_{\psi},(1+P)^{-\frac{s}{2}}](1+P)^{-\frac{\gamma}{2}}.$
By Theorem 3.1, the first factor belongs to $\mathcal{L}_{\frac{d}{\alpha+s+1},\infty}\cdot\mathcal{L}_{\infty}\cdot\mathcal{L}_{\frac{d}{\gamma},\infty}$ and therefore $\mathcal{L}_{\frac{d}{\alpha+\gamma+s+1},\infty}$. By Theorem 3.1, third factor belong to $\mathcal{L}_{\frac{d}{\alpha},\infty}\cdot\mathcal{L}_{\infty}\cdot\mathcal{L}_{\frac{d}{\gamma+s+1},\infty}$ and therefore $\mathcal{L}_{\frac{d}{\alpha+\gamma+s+1},\infty}$ Here $\mathcal{L}_{\frac{d}{\alpha},\infty}$ (or $\mathcal{L}_{\frac{d}{\gamma},\infty}$) is understood as $\mathcal{L}_{\infty}$ when $\alpha$ (or $\gamma$) is 0. For the second factor, Lemma 3.14 and Lemma 3.11 give that
$(1+P)^{-\frac{\alpha}{2}}M_{\psi}[M_{\phi}R_{j}M_{\phi},(1+P)^{-\frac{s}{2}}]M_{\psi}(1+P)^{-\frac{\gamma}{2}}$
$=(1+P)^{-\frac{\alpha}{2}}M_{\psi}(1+P)^{-\frac{1}{2}}\cdot(1+P)^{\frac{1-s}{2}}[(1+P)^{\frac{s}{2}},M_{\phi}R_{j}M_{\phi}]\cdot$
$\quad\cdot(1+P)^{-\frac{s}{2}}M_{\psi}(1+P)^{-\frac{\gamma}{2}}$
$\in\mathcal{L}_{\frac{d}{\alpha+1},\infty}\cdot\mathcal{L}_{\infty}\cdot\mathcal{L}_{\frac{d}{\gamma+s},\infty}\subset\mathcal{L}_{\frac{d}{\alpha+\gamma+s+1},\infty}.$
The proof is complete. ∎
4. Asymptotics on the curved plane
In this section, we generalize Theorem A to Theorem 4.1 by replacing the Laplacian with an elliptic differential operator on $\mathbb{R}^{d}$ for the index $p=\frac{d}{2}$.
Theorem 4.1.
Let $P:W^{2,2}(\mathbb{R}^{d})\to L_{2}(\mathbb{R}^{d})$ be a positive, self-adjoint, elliptic differential operator whose coefficients are constant outside some ball. If $S\in\Pi_{\mathbb{R}^{d}}$ is compactly supported, then
$\lim_{t\to\infty}t^{\frac{2}{d}}\mu(t,S(1+P)^{-1})=c_{d}\|{\rm sym}(S)\|_{L_{\frac{d}{2}}(\mathbb{R}^{d}\times\mathbb{R}^{d},e^{-\sigma_{P}}dm)}$
where $\sigma_{P}$ is the principal symbol of $P$.
Our starting point for proving Theorem 4.1 is the following lemma.
Lemma 4.2.
Let $P:W^{2,2}(\mathbb{R}^{d})\to L_{2}(\mathbb{R}^{d})$ be a positive, self-adjoint, elliptic differential operator whose coefficients are constant outside some ball. There exists $T\in\Pi_{\mathbb{R}^{d}}$ such that, for every $\phi\in C^{\infty}_{c}(\mathbb{R}^{d}),$
$(1+P)^{-1}M_{\phi}\in T(1+\Delta)^{-1}M_{\phi}+\mathcal{L}_{\frac{d}{3},\infty}(L_{2}(\mathbb{R}^{d}))$
and
${\rm sym}(T)(t,s)=(\sum_{i,j=1}^{d}a_{ij}(t)s_{i}s_{j})^{-1},\quad t\in\mathbb{R}^{d},\quad s\in\mathbb{S}^{d-1}$
where $\{a_{ij}\}_{i,j=1}^{d}$ are the coefficients of $P$ with respect to derivatives $\{D_{i}D_{j}\}_{i,j=1}^{d}$.
Proof.
Set
$\sigma(t,s)=\sum_{i,j=1}^{d}a_{ij}(t)s_{i}s_{j},\quad t,s\in\mathbb{R}^{d}.$
By the Complex Powers Theorem (see e.g. Theorem 1.6.22 in [24]), we have
$(1+P)^{-1}={\rm Op}((1+\sigma)^{-1})+P_{1},\quad P_{1}\in\Psi^{-3}(\mathbb{R}^{d}).$
Let $\psi\in C^{\infty}_{c}(\mathbb{R}^{d})$ be such that $\psi=1$ near $0.$ Set
$a(t,s)=(\sum_{i,j=1}^{d}a_{ij}(t)\frac{s_{i}}{|s|}\frac{s_{j}}{|s|})^{-1},\quad t,s\in\mathbb{R}^{d},$
and
$b(t,s)=a(t,s)\cdot(1-\psi)(s)\cdot(1+|s|^{2})^{-1},\quad t,s\in\mathbb{R}^{d}.$
Clearly,
$(1+\sigma)^{-1}-b\in S^{-3}(\mathbb{R}^{d}\times\mathbb{R}^{d}).$
Thus,
${\rm Op}((1+\sigma)^{-1})-{\rm Op}(b)\in\Psi^{-3}(\mathbb{R}^{d}).$
Using the preceding paragraph, we write
$(1+P)^{-1}={\rm Op}(b)+P_{2},\quad P_{2}\in\Psi^{-3}(\mathbb{R}^{d}).$
Next,
${\rm Op}(b)={\rm Op}(a)\cdot(1-\psi)(\nabla)\cdot(1+\Delta)^{-1}.$
Thus,
$(1+P)^{-1}={\rm Op}(a)\cdot(1-\psi)(\nabla)\cdot(1+\Delta)^{-1}+P_{2},\quad P_{2}\in\Psi^{-3}(\mathbb{R}^{d}).$
Clearly,
$P_{2}M_{\phi}\in\mathcal{L}_{\frac{d}{3},\infty}(L_{2}(\mathbb{R}^{d})).$
Hence,
$(1+P)^{-1}M_{\phi}\in{\rm Op}(a)\cdot(1-\psi)(\nabla)\cdot(1+\Delta)^{-1}M_{\phi}+\mathcal{L}_{\frac{d}{3},\infty}(L_{2}(\mathbb{R}^{d})).$
It remains to note that ${\rm Op}(a)\in\Pi_{\mathbb{R}^{d}}$ and its principal symbol
${\rm sym}({\rm Op}(a))(t,s)=(\sum_{i,j=1}^{d}a_{ij}(t)s_{i}s_{j})^{-1},\quad t\in\mathbb{R}^{d},\quad s\in\mathbb{S}^{d-1}.$
In addition,
$(1-\psi)(\nabla)\cdot(1+\Delta)^{-1}M_{\phi}\in(1+\Delta)^{-1}M_{\phi}+\mathcal{L}_{\frac{d}{3},\infty}(L_{2}(\mathbb{R}^{d})).$
This completes the proof. ∎
Lemma 4.3.
Let
$\sigma(t,s)=\sum_{i,j=1}^{d}a_{ij}(t)s_{i}s_{j},\quad t,s\in\mathbb{R}^{d}.$
For every $F\in L_{\infty}(\mathbb{R}^{d}\times\mathbb{S}^{d-1}),$ we have  [^1] Recall that we identify functions on the sphere with homogeneous functions on Euclidean space.
$\|F\cdot\sigma^{-1}\|_{L_{\frac{d}{2}}(\mathbb{R}^{d}\times\mathbb{S}^{d-1})}=\big(\frac{1}{2}\Gamma(\frac{d}{2})\big)^{-\frac{2}{d}}\|F\|_{L_{\frac{d}{2}}(\mathbb{R}^{d}\times\mathbb{R}^{d},e^{-\sigma}dm)}$
where $\Gamma(p):=\int_{0}^{\infty}x^{p-1}e^{-p}dx$, $p>0$, is the Gamma function.
Proof.
We have
$\|F\|_{L_{\frac{d}{2}}(\mathbb{R}^{d}\times\mathbb{R}^{d},e^{-\sigma}dm)}^{\frac{d}{2}}=\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}|F|^{\frac{d}{2}}(t,s)e^{-\sigma(t,s)}dtds.$
Let us pass to spherical coordinates $r=|s|,$ $u=\frac{s}{|s|}.$ We have
$|F|^{\frac{d}{2}}(t,s)=|F|^{\frac{d}{2}}(t,u),\quad e^{-\sigma(t,s)}=e^{-r^{2}\sigma(t,u)},\quad ds=r^{d-1}drdu.$
Thus,
$\|F\|_{L_{\frac{d}{2}}(\mathbb{R}^{d}\times\mathbb{R}^{d},e^{-\sigma}dm)}^{\frac{d}{2}}=\int_{\mathbb{R}^{d}\times\mathbb{S}^{d-1}}|F|^{\frac{d}{2}}(t,u)\Big(\int_{0}^{\infty}e^{-r^{2}\sigma(t,u)}r^{d-1}dr\Big)dtdu.$
Clearly,
$\int_{0}^{\infty}e^{-r^{2}\sigma(t,u)}r^{d-1}dr=\sigma(t,u)^{-\frac{d}{2}}\cdot\int_{0}^{\infty}e^{-x^{2}}x^{d-1}dx=\sigma(t,u)^{-\frac{d}{2}}\cdot\frac{1}{2}\Gamma(\frac{d}{2}).$
Finally,
$\|F\|_{L_{\frac{d}{2}}(\mathbb{R}^{d}\times\mathbb{R}^{d},e^{-\sigma}dm)}^{\frac{d}{2}}=\frac{1}{2}\Gamma(\frac{d}{2})\cdot\int_{\mathbb{R}^{d}\times\mathbb{S}^{d-1}}|F|^{\frac{d}{2}}(t,u)\sigma(t,u)^{-\frac{d}{2}}dtdu.$
This completes the proof. ∎
Now, we prove Theorem 4.1.
Proof of Theorem 4.1.
Observe that
$\lim_{t\to\infty}t^{\frac{2}{d}}\mu(t,S(1+P)^{-1})=\lim_{t\to\infty}t^{\frac{2}{d}}\mu(t,|S|(1+P)^{-1}),\quad S\in\Pi_{\mathbb{R}^{d}}.$
Without loss of generality, Suppose that $S\geq 0$.
Let $\phi\in C^{\infty}_{c}(\mathbb{R}^{d})$ be such that $S=M_{\phi}S.$ Let $T\in\Pi_{\mathbb{R}^{d}}$ be given by Lemma 4.2.
By Lemma 4.2, we have
$(1+P)^{-1}S=(1+P)^{-1}M_{\phi}S\in T(1+\Delta)^{-1}M_{\phi}S+\mathcal{L}_{\frac{d}{3},\infty}(L_{2}(\mathbb{R}^{d})).$
By [13, Theorem A.1],
$(1+\Delta)^{-1},M_{\phi}S]\in(\mathcal{L}_{\frac{d}{2},\infty})_{0}(L_{2}(\mathbb{R}^{d})).$
And, $\mathcal{L}_{\frac{d}{3},\infty}(L_{2}(\mathbb{R}^{d}))\subset(\mathcal{L}_{\frac{d}{2},\infty})_{0}(L_{2}(\mathbb{R}^{d}))$. Hence,
$(1+P)^{-1}S\in TM_{\phi}S(1+\Delta)^{-1}+(\mathcal{L}_{\frac{d}{2},\infty})_{0}(L_{2}(\mathbb{R}^{d})).$
Clearly, $T,M_{\phi}]$ is compact. Hence,
$(1+P)^{-1}S\in M_{\phi}TS(1+\Delta)^{-1}+(\mathcal{L}_{\frac{d}{2},\infty})_{0}(L_{2}(\mathbb{R}^{d})).$
The operator $M_{\phi}TS$ is compactly supported. By Theorem A, there exists a limit
$\lim_{t\rightarrow\infty}t^{\frac{2}{d}}\mu\Big(t,M_{\phi}TS(1+\Delta)^{-1}\Big)=c_{d}\|{\rm sym}(M_{\phi}TS)\|_{L_{\frac{d}{2}}(\mathbb{R}^{d}\times\mathbb{S}^{d-1})}.$
Clearly,
${\rm sym}(STM_{\phi})$ $={\rm sym}(S){\rm sym}(T){\rm sym}(M_{\phi})$
$={\rm sym}(S){\rm sym}(M_{\phi}){\rm sym}(T)$
$={\rm sym}(SM_{\phi}){\rm sym}(T)$
$={\rm sym}(S){\rm sym}(T).$
Hence,
$\lim_{t\rightarrow\infty}t^{\frac{2}{d}}\mu\Big(t,M_{\phi}TS(1+\Delta)^{-1}\Big)=c_{d}\|{\rm sym}(S){\rm sym}(T)\|_{L_{\frac{d}{2}}(\mathbb{R}^{d}\times\mathbb{S}^{d-1})}.$
The assertion follows now from Lemma 4.3. ∎
5. Powers of nearly commuting operators
In this section, we prove Theorem 5.1, the spectral characterization for powers of two nearly commutating operators, which substitutes the Complex Powers Theorem for pseudo-differential operators on manifold.
Theorem 5.1.
Let $(X,G)$ be a compact Riemannian manifold and let $\Delta_{G}$ be the Laplace-Beltrami operator. For every $0\leq S\in\Pi_{X},$ we have
$S^{s}(1+\Delta_{G})^{-s}-(S^{\frac{1}{2}}(1+\Delta_{G})^{-1}S^{\frac{1}{2}})^{s}\in(\mathcal{L}_{\frac{d}{2s},\infty})_{0}(L_{2}(X,{\rm vol}_{G})),\quad s>0.$
5.1. Abstract estimate
In this subsection, we prove Theorem 5.2. Firstly, fix $H$ a separable Hilbert space throughout this subsection.
Theorem 5.2.
Let $0\leq A\in\mathcal{L}_{\infty}(H)$ and let $0\leq B\in\mathcal{L}_{q,\infty}(H)$ with $q>0.$ If
$A^{\epsilon},B^{\delta}]\in(\mathcal{L}_{\frac{q}{\delta},\infty})_{0}(H)\quad\mbox{for any}\quad\epsilon,\delta>0,$
then
$B^{s}A^{s}-(A^{\frac{1}{2}}BA^{\frac{1}{2}})^{s}\in(\mathcal{L}_{\frac{q}{s},\infty})_{0}(H),\quad\mbox{if}\quad s>0.$
The following Lemma is a special case of Theorem 5.2 established in [8] (see Lemma 5.3 there).
Lemma 5.3.
Let $0\leq A\in\mathcal{L}_{\infty}(H)$ and let $0\leq B\in\mathcal{L}_{p,\infty}(H)$ with $p>1.$ If $A^{\frac{1}{2}},B]\in(\mathcal{L}_{p,\infty})_{0}(H),$ then
$B^{p}A^{p}-(A^{\frac{1}{2}}BA^{\frac{1}{2}})^{p}\in(\mathcal{L}_{1,\infty})_{0}(H).$
Lemma 5.4.
If $0\leq X,Y\in\mathcal{L}_{r,\infty}(H)$ and $X-Y\in(\mathcal{L}_{r,\infty})_{0}(H)$ with $r>0$, then
$X^{t}-Y^{t}\in(\mathcal{L}_{\frac{r}{t},\infty})_{0}(H),\quad t>0.$
Proof.
Fix $n\in\mathbb{N}$ such that $n>t.$ Since $X-Y\in(\mathcal{L}_{r,\infty})_{0}(H),$ applying the proof of [24, Corollary 1.3.15] and the most general form of the Birman-Koplienko-Solomyak inequality established in [18], we write
$X^{\frac{t}{n}}-Y^{\frac{t}{n}}\in(\mathcal{L}_{\frac{nr}{t},\infty})_{0}(H).$
Now, we write
$X^{t}-Y^{t}=\sum_{k=0}^{n-1}X^{\frac{kt}{n}}(X^{\frac{t}{n}}-Y^{\frac{t}{n}})Y^{\frac{(n-1-k)t}{n}}.$
The assertion follows now from Hölder inequality. ∎
The next lemma is an easy application of Lemma 5.3 and Lemma 5.4.
Proof of Theorem 5.2.
Choose $n\in\mathbb{N}$ such that $ns>1.$ Choose $m\in\mathbb{N}$ such that $mq>ns$ and $mq>1.$
By assumption, $0\leq A^{\frac{ns}{m}}\in\mathcal{L}_{\infty}(H)$ and $0\leq B^{\frac{ns}{m}}\in\mathcal{L}_{\frac{mq}{ns},\infty}(H)$ are such that
$A^{\frac{ns}{2m}},B^{\frac{ns}{m}}]\in(\mathcal{L}_{\frac{mq}{ns},\infty})_{0}(H).$
Applying Lemma 5.3 with $p=\frac{mq}{ns}$ to the couple $(A^{\frac{ns}{m}},B^{\frac{ns}{m}}),$ we obtain
(5.1) $B^{q}A^{q}-(A^{\frac{ns}{2m}}B^{\frac{ns}{m}}A^{\frac{ns}{2m}})^{\frac{mq}{ns}}\in(\mathcal{L}_{1,\infty})_{0}(H).$
By assumption, $0\leq A^{\frac{n}{m}}\in\mathcal{L}_{\infty}(H)$ and $0\leq B^{\frac{n}{m}}\in\mathcal{L}_{\frac{mq}{n},\infty}(H)$ are such that
$A^{\frac{n}{2m}},B^{\frac{n}{m}}]\in(\mathcal{L}_{\frac{mq}{n},\infty})_{0}(H).$
Applying Lemma 5.3 with $p=\frac{mq}{n}$ to the couple $(A^{\frac{n}{m}},B^{\frac{n}{m}}),$ we obtain
(5.2) $B^{q}A^{q}-(A^{\frac{n}{2m}}B^{\frac{n}{m}}A^{\frac{n}{2m}})^{\frac{mq}{n}}\in(\mathcal{L}_{1,\infty})_{0}(H).$
Combining (5.1) and (5.2), we write
$(A^{\frac{ns}{2m}}B^{\frac{ns}{m}}A^{\frac{ns}{2m}})^{\frac{mq}{ns}}-(A^{\frac{n}{2m}}B^{\frac{n}{m}}A^{\frac{n}{2m}})^{\frac{mq}{n}}\in(\mathcal{L}_{1,\infty})_{0}(H).$
Using Lemma 5.4 with
$X=(A^{\frac{ns}{2m}}B^{\frac{ns}{m}}A^{\frac{ns}{2m}})^{\frac{mq}{ns}},\quad Y=(A^{\frac{n}{2m}}B^{\frac{n}{m}}A^{\frac{n}{2m}})^{\frac{mq}{n}},\quad r=1,\quad t=\frac{ns}{mq},$
we write
(5.3) $A^{\frac{ns}{2m}}B^{\frac{ns}{m}}A^{\frac{ns}{2m}}-(A^{\frac{n}{2m}}B^{\frac{n}{m}}A^{\frac{n}{2m}})^{s}\in(\mathcal{L}_{\frac{mq}{ns},\infty})_{0}(H).$
By assumption,
$A^{\frac{n}{2m}}B^{\frac{n}{m}}A^{\frac{n}{2m}}-(A^{\frac{1}{2m}}B^{\frac{1}{m}}A^{\frac{1}{2m}})^{n}\in(\mathcal{L}_{\frac{mq}{n},\infty})_{0}(H).$
Using Lemma 5.4 with
$X=A^{\frac{n}{2m}}B^{\frac{n}{m}}A^{\frac{n}{2m}},\quad Y=(A^{\frac{1}{2m}}B^{\frac{1}{m}}A^{\frac{1}{2m}})^{n},\quad r=\frac{mq}{n},\quad t=s,$
we write
(5.4) $(A^{\frac{n}{2m}}B^{\frac{n}{m}}A^{\frac{n}{2m}})^{s}-(A^{\frac{1}{2m}}B^{\frac{1}{m}}A^{\frac{1}{2m}})^{ns}\in(\mathcal{L}_{\frac{mq}{ns},\infty})_{0}(H).$
By assumption,
(5.5) $A^{\frac{ns}{2m}}B^{\frac{ns}{m}}A^{\frac{ns}{2m}}-(A^{\frac{s}{2m}}B^{\frac{s}{m}}A^{\frac{s}{2m}})^{n}\in(\mathcal{L}_{\frac{mq}{ns},\infty})_{0}(H).$
Combining (5.3), (5.4) and (5.5), we write
$(A^{\frac{s}{2m}}B^{\frac{s}{m}}A^{\frac{s}{2m}})^{n}-(A^{\frac{1}{2m}}B^{\frac{1}{m}}A^{\frac{1}{2m}})^{ns}\in(\mathcal{L}_{\frac{mq}{ns},\infty})_{0}(H).$
Using Lemma 5.4 with
$X=(A^{\frac{s}{2m}}B^{\frac{s}{m}}A^{\frac{s}{2m}})^{n},\quad Y=(A^{\frac{1}{2m}}B^{\frac{1}{m}}A^{\frac{1}{2m}})^{ns},\quad r=\frac{mq}{ns},\quad s=\frac{1}{n},$
we write
(5.6) $A^{\frac{s}{2m}}B^{\frac{s}{m}}A^{\frac{s}{2m}}-(A^{\frac{1}{2m}}B^{\frac{1}{m}}A^{\frac{1}{2m}})^{s}\in(\mathcal{L}_{\frac{mq}{s},\infty})_{0}(H).$
By assumption,
$(A^{\frac{1}{2m}}B^{\frac{1}{m}}A^{\frac{1}{2m}})^{m}-A^{\frac{1}{2}}BA^{\frac{1}{2}}\in(\mathcal{L}_{q,\infty})_{0}(H).$
Using Lemma 5.4 with
$X=(A^{\frac{1}{2m}}B^{\frac{1}{m}}A^{\frac{1}{2m}})^{m},\quad Y=A^{\frac{1}{2}}BA^{\frac{1}{2}},\quad r=q,\quad t=\frac{s}{m},$
we write
(5.7) $(A^{\frac{1}{2m}}B^{\frac{1}{m}}A^{\frac{1}{2m}})^{s}-(A^{\frac{1}{2}}BA^{\frac{1}{2}})^{\frac{s}{m}}\in(\mathcal{L}_{\frac{mq}{s},\infty})_{0}(H).$
Combining (5.6) and (5.7), we write
$A^{\frac{s}{2m}}B^{\frac{s}{m}}A^{\frac{s}{2m}}-(A^{\frac{1}{2}}BA^{\frac{1}{2}})^{\frac{s}{m}}\in(\mathcal{L}_{\frac{mq}{s},\infty})_{0}(H).$
By Lemma 5.4,
$(A^{\frac{s}{2m}}B^{\frac{s}{m}}A^{\frac{s}{2m}})^{m}-(A^{\frac{1}{2}}BA^{\frac{1}{2}})^{s}\in(\mathcal{L}_{\frac{q}{s},\infty})_{0}(H).$
By assumption,
$(A^{\frac{s}{2m}}B^{\frac{s}{m}}A^{\frac{s}{2m}})^{m}-A^{\frac{s}{2}}B^{s}A^{\frac{s}{2}}\in(\mathcal{L}_{\frac{q}{s},\infty})_{0}(H).$
Combining the last two inclusions, we complete the proof. ∎
5.2. Application of abstract estimate
We prove Theorem 5.1 in this subsection. To do that, we introduce the notion for differentiability of operators inspiring by Connes–Moscovici [3, page 209]. See the following definition on $\mathbb{R}^{d}$.
Definition 5.5.
Element $S\in\Pi_{\mathbb{R}^{d}}$ is called $m$ times differentiable if, for every second order differential operator $P$ with smooth compactly supported coefficients, $k$-th commutator
$\underbrace{[P,[P,\cdots[P,S]\cdots]}_{k\text{ times}}$
extends to a bounded operator from $W^{k,2}(\mathbb{R}^{d})\to L_{2}(\mathbb{R}^{d})$ for every $1\leq k\leq m.$
The similar notion on compact manifold can be proposed in the following way.
Definition 5.6.
Let $(X,G)$ be a compact manifold. Element $S\in\Pi_{X}$ is called $m$ times differentiable if, for every second order differential operator $P$ with smooth coefficients, $k$-th commutator
$\underbrace{[P,[P,\cdots[P,S]\cdots]}_{k\text{ times}}$
extends to a bounded operator from $W^{k,2}(X,{\rm vol}_{G})\to L_{2}(X,{\rm vol}_{G})$ for every $1\leq k\leq m.$
Let us turn to the density of differentiable elements. Notably, the algebra $\{g(\frac{\nabla}{\sqrt{\Delta}}):g\in C(\mathbb{S}^{d-1})\}$ can also be generated by the Riesz transforms $R_{j}=\partial_{j}\Delta^{-\frac{1}{2}}$, $j=1,\cdots,d$. We have below results.
Lemma 5.7.
Twice differentiable elements are dense in $\Pi_{\mathbb{R}^{d}}$.
Proof.
Let $P$ be a second order differential operator with smooth compactly supported coefficients. Let us firstly claim that $\{R_{j}\}_{j=1}^{d}$ are twice differentiable. Clearly, $P\in\Psi^{2}(\mathbb{R}^{d})$ and $R_{j}\in\Psi^{0}(\mathbb{R}^{d})$ for $j=1,\cdots,d$. By gradation of pseudo-differentiable operators,
$P,R_{j}](1+\Delta)^{-\frac{1}{2}},\quad[P,[P,R_{j}]](1+\Delta)^{-1}\in\Psi^{0}(\mathbb{R}^{d})$
where we use the fact $(1+\Delta)^{-\frac{1}{2}}\in\Psi^{-1}(\mathbb{R}^{d})$ and $(1+\Delta)^{-1}\in\Psi^{-2}(\mathbb{R}^{d})$. The boundedness on $L_{2}(\mathbb{R}^{d})$ of these two commutators follows from [43, Theorem 2,page 286] and then the claim follows.
Similarly, the operator $M_{f}$ with $f\in C_{c}^{\infty}(\mathbb{R}^{d})$ is twice differentiable since $M_{f}\in\Psi^{0}(\mathbb{R}^{d})$.
Let $S\in\Pi_{\mathbb{R}^{d}}$. By the definition of $\Pi_{\mathbb{R}^{d}},$ we can find a sequence $\{S_{n}\}_{n\geq 0}$ such that each $S_{n}$ belongs to $\ast$-algebra generated by $\{M_{f}:f\in C_{0}(\mathbb{R}^{d})\}$ and by $\{R_{j}\}_{j=1}^{d}$, and such that $S_{n}\to S$ in the uniform norm. Since $C_{c}^{\infty}(\mathbb{R}^{d})$ is dense in $C_{0}(\mathbb{R}^{d})$ in the uniform norm, there is a sub-sequence $\{S_{n_{k}}\}_{k\geq 0}$ generated by $\{M_{f}:f\in C_{c}^{\infty}(\mathbb{R}^{d})\}$ and $\{R_{j}\}_{j=1}^{d}$, and such that $S_{n_{k}}\to S$ in the uniform norm. By arguments in the preceding two paragraphs, $S_{n_{k}}$ is twice differentiable. The proof is complete. ∎
Lemma 5.8.
Let $(X,G)$ be a compact Riemannian manifold. Twice differentiable elements compactly supported in the chart $({\mathcal{U}}_{i},h_{i})$ are dense in the set of elements of $\Pi_{X}$ compactly supported in the chart $({\mathcal{U}}_{i},h_{i}).$
Proof.
Let $(\mathcal{U}_{i},h_{i})$ be a chart and let $S\in\Pi_{X}$ be compactly supported in $\mathcal{U}_{i}.$ The operator $W_{h_{i}}SW_{h_{i}}^{\ast}\in\Pi_{\mathbb{R}^{d}}$ is compactly supported in $h_{i}(\mathcal{U}_{i}).$ Fix $\phi_{i}\in C^{\infty}_{c}(h_{i}(\mathcal{U}_{i}))$ such that
$W_{h_{i}}SW_{h_{i}}^{\ast}=W_{h_{i}}SW_{h_{i}}^{\ast}\cdot M_{\phi_{i}}=M_{\phi_{i}}\cdot W_{h_{i}}SW_{h_{i}}^{\ast}.$
By Lemma 5.7, there exists a sequence $\{S_{n}\}_{n\geq 1}\subset\Pi_{\mathbb{R}^{d}}$ such that (i) $S_{n}$ is twice differentiable for every $n\geq 1$ and (ii) $S_{n}\to W_{h_{i}}SW_{h_{i}}^{\ast}$ in the uniform norm. Since $M_{\phi_{i}}$ is [^2] This is established in the proof of Lemma 5.7. twice differentiable, it follows that so is $M_{\phi_{i}}S_{n}M_{\phi_{i}}$ for every $n\geq 1.$ We have
$M_{\phi_{i}}S_{n}M_{\phi_{i}}\to M_{\phi_{i}}\cdot W_{h_{i}}SW_{h_{i}}^{\ast}\cdot M_{\phi_{i}}=W_{h_{i}}SW_{h_{i}}^{\ast}$
in the uniform norm. Replacing $S_{n}$ with $M_{\phi_{i}}S_{n}M_{\phi_{i}}$ if needed, we may assume without loss of generality that each $S_{n}$ is compactly supported in $h_{i}(\mathcal{U}_{i}).$
The operator $W_{h_{i}}^{\ast}S_{n}W_{h_{i}}\in\Pi_{X}$ is twice differentiable for every $n\geq 1.$ Clearly, $W_{h_{i}}^{\ast}S_{n}W_{h_{i}}\to S$ in the uniform norm. This completes the proof. ∎
Lemma 5.9.
Let $(X,G)$ be a compact Riemannian manifold. Twice differentiable elements are dense in $\Pi_{X}$ in the uniform norm.
Proof.
Let $\{\phi_{n}\}_{n=1}^{N}\subset C_{c}^{\infty}(X)$ be a partition of unity subordinate to the atlas of $X.$ For $S\in\Pi_{X},$ we write
$S=\sum_{n=1}^{N}M_{\phi_{n}^{\frac{1}{2}}}SM_{\phi_{n}^{\frac{1}{2}}}+\sum_{n=1}^{N}[S,M_{\phi_{n}^{\frac{1}{2}}}]M_{\phi_{n}^{\frac{1}{2}}}.$
The approximation for summands in the first sum on the right hand side are given by Lemma 5.8. Observe that the summands in the second sum on the right hand side are compact operators. Since integral operators with smooth integral kernels are differentiable, compact and dense among compact operators in the uniform norm, the assertion follows. ∎
Lemma 5.10.
Let $(X,G)$ be a compact Riemannian manifold. If $S\in\Pi_{X}$ is twice differentiable, then $S,(1+\Delta_{G})^{\frac{1}{2}}]$ is bounded.
Proof.
Since $X$ is compact, the assumption can be equivalently restated as the boundedness of operators
$\Delta_{G},S](1+\Delta_{G})^{-\frac{1}{2}},[\Delta_{G},[\Delta_{G},S]](1+\Delta_{G})^{-1}$
on $L_{2}(X,{\rm vol}_{G}).$ Recall that
$(1+\Delta_{G})^{-\frac{1}{2}}=\frac{1}{\pi}\int_{0}^{\infty}(1+\lambda+\Delta_{G})^{-1}\lambda^{-\frac{1}{2}}d\lambda$
where the convergence for the integral is understood in the operator norm. Write
(5.8) $[(1+\Delta_{G})^{\frac{1}{2}},S]=[(1+\Delta_{G})(1+\Delta_{G})^{-\frac{1}{2}},S]=\frac{1}{\pi}\int_{0}^{\infty}[\frac{1+\Delta_{G}}{1+\lambda+\Delta_{G}},S]\lambda^{-\frac{1}{2}}d\lambda.$
However,
$[\frac{1+\Delta_{G}}{1+\lambda+\Delta_{G}},S]$ $=[\Delta_{G},S](1+\Delta_{G})^{-\frac{1}{2}}\cdot\Big(\frac{(1+\Delta_{G})^{\frac{1}{2}}}{1+\lambda+\Delta_{G}}-\frac{(1+\Delta_{G})^{\frac{3}{2}}}{(1+\lambda+\Delta_{G})^{2}}\Big)+$
$+\frac{\lambda}{1+\lambda+\Delta_{G}}\cdot[\Delta_{G},[\Delta_{G},S]](1+\Delta_{G})^{-1}\cdot\frac{1+\Delta_{G}}{(1+\lambda+\Delta_{G})^{2}}$
$=:X_{\lambda}+Y_{\lambda}.$
By functional calculus,
$\int_{0}^{\infty}\Big(\frac{(1+\Delta_{G})^{\frac{1}{2}}}{1+\lambda+\Delta_{G}}-\frac{(1+\Delta_{G})^{\frac{3}{2}}}{(1+\lambda+\Delta_{G})^{2}}\Big)\lambda^{-\frac{1}{2}}d\lambda=\frac{\pi}{2}.$
It follows that
$\frac{1}{\pi}\int_{0}^{\infty}X_{\lambda}\cdot\lambda^{-\frac{1}{2}}d\lambda=\frac{1}{2}[\Delta_{G},S](1+\Delta_{G})^{-\frac{1}{2}}$
is bounded on $L_{2}(X,{\rm vol}_{G})$ by the assumption. Additionally, by spectral theorem,
$\|\frac{\lambda}{1+\lambda+\Delta_{G}}\|_{B(L_{2}(X,{\rm vol}_{G}))}=\frac{\lambda}{1+\lambda}$
and
$\|\frac{1+\Delta_{G}}{(1+\lambda+\Delta_{G})^{2}}\|_{B(L_{2}(X,{\rm vol}_{G}))}=\left\{\begin{array}[]{cl}\frac{1}{(1+\lambda)^{2}},\quad&0<\lambda\leq 1,\\ \frac{1}{4\lambda},\quad&\lambda\geq 1.\end{array}\right$.
Thus,
$\|\int_{0}^{\infty}Y_{\lambda}\cdot\lambda^{-\frac{1}{2}}d\lambda\|_{B(L_{2}(X,{\rm vol}_{G}))}\leq\frac{7}{6}\|[\Delta_{G},[\Delta_{G},S]](1+\Delta_{G})^{-1}\|_{B(L_{2}(X,{\rm vol}_{G}))}.$
Taking the arguments for $X_{\lambda},Y_{\lambda}$, equality (5.8) and the assumption for $S$ into a combination, the assertion follows. ∎
Next, we verify the assumptions of Theorem 5.2 on manifold $(X,G)$.
Lemma 5.11.
Let $(X,G)$ be a compact Riemannian manifold and let $\Delta_{G}$ be the Laplace-Beltrami operator. If $0\leq S\in\Pi_{X},$ then $A=S$ and $B=(1+\Delta_{G})^{-1}$ satisfy the assumptions in Theorem 5.2 with $q=\frac{d}{2}.$
Proof.
Let $\{S_{n}\}_{n\geq 1}\subset\Pi_{X}$ be a sequence of twice differentiable elements such that $S_{n}\to S^{\epsilon}$ in the uniform norm.
Fix $m>\delta$ and write
$[S_{n},B^{\delta}]=[S_{n},(B^{\frac{\delta}{2m}})^{2m}]$ $=\sum_{k=0}^{2m-1}(B^{\frac{\delta}{2m}})^{k}\cdot[S_{n},B^{\frac{\delta}{2m}}]\cdot(B^{\frac{\delta}{2m}})^{2m-1-k}$
$=-\sum_{k=0}^{2m-1}(B^{\frac{\delta}{2m}})^{k+1}\cdot[S_{n},B^{-\frac{\delta}{2m}}]\cdot(B^{\frac{\delta}{2m}})^{2m-k}.$
As $S_{n}$ is twice differentiable, the operator $S_{n},B^{-\frac{1}{2}}]$ is bounded due to Lemma 5.10. By Birman-Koplienko-Solomyak inequality [24, Theorem 1.3.4],
$\int_{0}^{t}\mu(s,[S_{n},B^{-\frac{\delta}{2m}}])ds\leq c_{\delta,m}\|S_{n}\|_{B(L_{2}(X,{\rm vol}_{G}))}^{1-\frac{\delta}{m}}\int_{0}^{t}\mu(s,|[S_{n},B^{-\frac{1}{2}}]|^{\frac{\delta}{m}})ds,\quad t>0.$
Thus, the commutator $S_{n},B^{-\frac{\delta}{2m}}]$ is bounded. It follows now from Hölder inequality that
$S_{n},B^{\delta}]\in\mathcal{L}_{\frac{dm}{(2m+1)\delta},\infty}(L_{2}(X,{\rm vol}_{G}))\subset(\mathcal{L}_{\frac{d}{2\delta}})_{0}(L_{2}(X,{\rm vol}_{G})).$
Since $S_{n},B^{\delta}]\to[A^{\epsilon},B^{\delta}]\in\mathcal{L}_{\frac{d}{2\delta}}(L_{2}(X,{\rm vol}_{G}))$ as $n\to\infty,$ the assertion follows from Lemma 5.9. The proof is complete. ∎
Proof of Theorem 5.1.
It is established in Lemma 5.11 that the operators
$A=S,\quad B=(1+\Delta_{G})^{-1}$
satisfy the assumptions in Theorem 5.2 with $q=\frac{d}{2}.$ Applying Theorem 5.2, we complete the proof. ∎
6. Asymptotics for operators localised in a chart
This section is devoted to the proof of Theorem 6.7, which establishes a local version of the result stated in Theorem 1.2.
6.1. Good atlas on compact Riemannian manifold
To show Theorem 6.7, we introduce the definition of good atlas on compact manifold.
Definition 6.1.
Let $(X,G)$ be a compact Riemannian manifold. A finite atlas $\{({\mathcal{U}}_{i},h_{i})\}_{i\in\mathbb{I}}$ is called good if, for every $i\in\mathbb{I},$ the function $g_{i}:h_{i}({\mathcal{U}}_{i})\to{\rm GL}_{+}(d,\mathbb{R})$ allows a smooth extension $g_{i}:\mathbb{R}^{d}\to{\rm GL}_{+}(d,\mathbb{R})$ which is constant outside some ball.
Towards constructing good atlas on compact Riemannian manifolds, the following smoothness result is established as a start.
Lemma 6.2.
Let $\Omega$ be a domain in $\mathbb{R}^{d}$ and let $g:\Omega\to{\rm GL}_{+}(d,\mathbb{R})$ be a smooth function. The function $\log\circ g:\Omega\to M_{d}(\mathbb{R})$ is also smooth where $M_{d}(\mathbb{R})$ denotes the space of all $d\times d$ real matrices.
Proof.
Fix $t_{0}\in\Omega.$ Since $g(t_{0})$ is a positive matrix, its spectrum is positive. Let $\lambda_{\text{min}}(t_{0})$ and $\lambda_{\text{max}}(t_{0})$ denote its smallest and largest eigenvalues, respectively. We have
$\|g(t_{0})\|_{M_{d}(\mathbb{R})}=\lambda_{\text{max}}(t_{0}),\quad\|g(t_{0})^{-1}\|_{M_{d}(\mathbb{R})}=\frac{1}{\lambda_{\text{min}}(t_{0})},$
where $\|\cdot\|_{M_{d}(\mathbb{R})}$ denotes the operator norm on $M_{d}(\mathbb{R})$. Therefore,
$\|g(t_{0})^{-1}\|_{M_{d}(\mathbb{R})}^{-1}\leq g(t_{0})\leq\|g(t_{0})\|_{M_{d}(\mathbb{R})}.$
Choose $\epsilon>0$ such that
$\|g(t)-g(t_{0})\|_{M_{d}(\mathbb{R})}<\frac{1}{2}\|g(t_{0})^{-1}\|_{M_{d}(\mathbb{R})}^{-1},\quad|t-t_{0}|<\epsilon.$
We have
$\frac{1}{2}\|g(t_{0})^{-1}\|_{M_{d}(\mathbb{R})}^{-1}\leq g(t)\leq\frac{1}{2}\|g(t_{0})^{-1}\|_{M_{d}(\mathbb{R})}^{-1}+\|g(t_{0})\|_{M_{d}(\mathbb{R})},\quad|t-t_{0}|<\epsilon.$
Choose $\omega\in C^{\infty}(\mathbb{R})$ such that
$\omega(u)=\log(u),\quad u\in[\frac{1}{2}\|g(t_{0})^{-1}\|_{M_{d}(\mathbb{R})}^{-1},\frac{1}{2}\|g(t_{0})^{-1}\|_{M_{d}(\mathbb{R})}^{-1}+\|g(t_{0})\|_{M_{d}(\mathbb{R})}].$
The mapping $\omega\circ g$ is obviously smooth on $\Omega$. Note that $\log\circ g=\omega\circ g$ on $B(t_{0},\epsilon).$ Hence, $\log\circ g$ is smooth on $B(t_{0},\epsilon).$ Since $t_{0}\in\Omega$ is arbitrary, the assertion follows. ∎
A compact manifold admits a finite atlas. This allows us to derive a more refined atlas in the setting of compact Riemannian manifolds.
Proposition 6.3.
Every compact Riemannian manifold allows a good atlas.
Proof.
Let $(X,G)$ be a compact Riemannian manifold with atlas $\{({\mathcal{U}}_{i},h_{i})\}_{i\in\mathbb{I}}.$ For every open set $D$ compactly supported in $h_{i}({\mathcal{U}}_{i})$ (i.e. the closure $\overline{D}$ is compact and $\overline{D}\subset h_{i}(\mathcal{U}_{i})$), choose a function $\phi_{D}\in C^{\infty}_{c}(h_{i}({\mathcal{U}}_{i}))$ such that $\phi_{D}=1$ on $D.$ Since $X$ is a Riemannian manifold, the function $g_{i}:h_{i}(\mathcal{U}_{i})\to{\rm GL}_{+}(d,\mathbb{R})$ is well-defined and smooth. By Lemma 6.2, the function $\log\circ g_{i}:h_{i}({\mathcal{U}}_{i})\to M_{d}(\mathbb{R})$ is smooth on $h_{i}({\mathcal{U}}_{i}).$ Hence, $g_{i}:D\to{\rm GL}_{+}(d,\mathbb{R})$ allows an extension $g_{i,D}:\mathbb{R}^{d}\to{\rm GL}_{+}(d,\mathbb{R})$ given by the formula
$g_{i,D}:x\to\exp(\phi_{D}(x)\cdot\log(g_{i}(x))),\quad x\in\mathbb{R}^{d}.$
Clearly, $g_{i,D}$ is a smooth extension of $g_{i}|_{D}$ and it is constant outside some ball.
Consider now the new atlas  [^3] Here, $D\Subset h_{i}({\mathcal{U}}_{i})$ means ”$D$ is compactly supported in $h_{i}({\mathcal{U}}_{i})$”. $\{\{(h_{i}^{-1}(D),h_{i})\}_{D\Subset h_{i}({\mathcal{U}}_{i})}\}_{i\in\mathbb{I}}.$ In this atlas, Riemannian metric $V$ is given by the formula $V_{(i,D)}=G_{i}|_{h_{i}^{-1}(D)}$ and the coordinate mapping $h_{(i,D)}$ is $h_{i}$ related to $h_{i}^{-1}(D)$. For every index $(i,D),$ we have $V_{(i,D)}\circ h_{(i,D)}^{-1}=g_{i,D}|_{D}=g_{i}|_{D}.$ Finally, by compactness on $X$, there exists a finite sub-atlas and it is good. The proof is complete. ∎
6.2. Local case for $p=\frac{d}{2}$
This subsection aims to show Theorem 6.7. We need the necessary lemmas from below.
Lemma 6.4.
Let $(X,G)$ be a compact Riemannian manifold and let $\Delta_{G}$ be the Laplace-Beltrami operator. Let $({\mathcal{U}}_{i},h_{i})$ be a chart of $X$. For $\phi\in C^{\infty}_{c}({\mathcal{U}}_{i}),$ the operator $\Delta_{G},M_{\phi}](1+\Delta_{G})^{-\frac{1}{2}}$ is bounded.
Proof.
Clearly, $M_{\phi}\in\Psi^{0}(X)$, $\Delta_{G}\in\Psi^{2}(X)$ and $(1+\Delta_{G})^{-\frac{1}{2}}\in\Psi^{-1}(X)$. We have $\Delta_{G},M_{\phi}]\in\Psi^{1}(X)$ by [34, Remark 5.3.3]. The composition of pseudo-differential operators gives that $\Delta_{G},M_{\phi}](1+\Delta_{G})^{-\frac{1}{2}}\in\Psi^{0}(X).$ Since $X$ is a compact manifold, the assertion follows from [34, Theorem 5.2.23]. ∎
Lemma 6.5.
Let $(X,G)$ be a compact Riemannian manifold and let $\Delta_{G}$ be the Laplace-Beltrami operator. Let $({\mathcal{U}}_{i},h_{i})$ be a chart in a good (in the sense of Definition 6.1) atlas. For every $\phi\in C^{\infty}_{c}({\mathcal{U}}_{i}),$ we have
$M_{\phi}(1+\Delta_{G})^{-1}M_{\phi}-W_{h_{i}}^{\ast}M_{\phi\circ h_{i}^{-1}}(1+\Delta_{g_{i}})^{-1}M_{\phi\circ h_{i}^{-1}}W_{h_{i}}\in\mathcal{L}_{\frac{d}{3},\infty}(L_{2}(X,{\rm vol}_{G})).$
Proof.
We prove an equivalent assertion
$W_{h_{i}}M_{\phi}(1+\Delta_{G})^{-1}M_{\phi}W_{h_{i}^{\ast}}-M_{\phi\circ h_{i}^{-1}}(1+\Delta_{g_{i}})^{-1}M_{\phi\circ h_{i}^{-1}}\in\mathcal{L}_{\frac{d}{3},\infty}(L_{2}(\mathbb{R}^{d},{\rm vol}_{g_{i}})).$
Denote the operator on the left hand side in the latter inclusion by $LHS.$
Note that $LHS$ sends $L_{2}(\mathbb{R}^{d},{\rm vol}_{g_{i}})$ into $W^{2,2}(\mathbb{R}^{d},{\rm vol}_{g_{i}}),$ which is the domain of $\Delta_{g_{i}}.$ We have
$(1+\Delta_{g_{i}})\cdot LHS=$
$=W_{h_{i}}(1+\Delta_{G})M_{\phi}(1+\Delta_{G})^{-1}M_{\phi}W_{h_{i}^{\ast}}-(1+\Delta_{g_{i}})M_{\phi\circ h_{i}^{-1}}(1+\Delta_{g_{i}})^{-1}M_{\phi\circ h_{i}^{-1}}$
$=W_{h_{i}}[\Delta_{G},M_{\phi}](1+\Delta_{G})^{-1}M_{\phi}W_{h_{i}^{\ast}}-[\Delta_{g_{i}},M_{\phi\circ h_{i}^{-1}}](1+\Delta_{g_{i}})^{-1}M_{\phi\circ h_{i}^{-1}}.$
In other words,
$LHS=(1+\Delta_{g_{i}})^{-1}\cdot\Big(W_{h_{i}}ABW_{h_{i}}^{\ast}-CD\Big),$
where
$A=[\Delta_{G},M_{\phi}](1+\Delta_{G})^{-\frac{1}{2}},\quad B=(1+\Delta_{G})^{-\frac{1}{2}}M_{\phi},$
$C=[\Delta_{g_{i}},M_{\phi\circ h_{i}^{-1}}](1+\Delta_{g_{i}})^{-\frac{1}{2}},\quad D=(1+\Delta_{g_{i}})^{-\frac{1}{2}}M_{\phi\circ h_{i}^{-1}}.$
Choose $\psi\in C^{\infty}_{c}(\mathbb{R}^{d})$ such that $\psi=1$ on and near ${\rm supp}(\phi\circ h_{i}^{-1}).$ It follows that
(6.1) $LHS=(1+\Delta_{g_{i}})^{-1}M_{\psi}\cdot\Big(W_{h_{i}}ABW_{h_{i}}^{\ast}-CD\Big).$
By Lemma 6.4 and Lemma 2.3,
$A\in\mathcal{L}_{\infty}(L_{2}(X,{\rm vol}_{G})),\quad B\in\mathcal{L}_{d,\infty}(L_{2}(X,{\rm vol}_{G})).$
Hence,
(6.2) $W_{h_{i}}ABW_{h_{i}}^{\ast}\in\mathcal{L}_{d,\infty}(L_{2}(X,{\rm vol}_{G})).$
On the other hand, consider $M_{{\rm det}^{\frac{1}{4}}(g_{i})}:L_{2}(\mathbb{R}^{d},{\rm vol}_{g_{i}})\to L_{2}(\mathbb{R}^{d})$ the unitary mapping. We have
$M_{{\rm det}^{-\frac{1}{4}}(g_{i})}[\Delta_{g_{i}},M_{\phi\circ h_{i}^{-1}}](1+\Delta_{g_{i}})^{-\frac{1}{2}}M_{{\rm det}^{\frac{1}{4}}(g_{i})}=[P_{i},M_{\phi\circ h_{i}^{-1}}](1+P_{i})^{-\frac{1}{2}}$
where $P_{i}=M_{{\rm det}^{-\frac{1}{4}}(g_{i})}\Delta_{g_{i}}M_{{\rm det}^{\frac{1}{4}}(g_{i})}$ is a positive, self-adjoint, elliptic differential operator whose coefficients are smooth and also are constant outside some ball on $\mathbb{R}^{d}$. By Lemma 3.8,
$P_{i},M_{\phi\circ h_{i}^{-1}}](1+P_{i})^{-\frac{1}{2}}\in\mathcal{L}_{\infty}(L_{2}(\mathbb{R}^{d}))$
and therefore
$C\in\mathcal{L}_{\infty}(L_{2}(\mathbb{R}^{d},{\rm vol}_{g_{i}})).$
Moreover, write
$M_{{\rm det}^{-\frac{1}{4}}(g_{i})}(1+\Delta_{g_{i}})^{-\frac{1}{2}}M_{\phi\circ h_{i}^{-1}}M_{{\rm det}^{\frac{1}{4}}(g_{i})}=(1+P_{i})^{-\frac{1}{2}}M_{\phi\circ h_{i}^{-1}}.$
The Lemma 3.11 gives that
$(1+P_{i})^{-\frac{1}{2}}M_{\phi\circ h_{i}^{-1}}\in\mathcal{L}_{d,\infty}(L_{2}(\mathbb{R}^{d}))$
and therefore
$D\in\mathcal{L}_{d,\infty}(L_{2}(\mathbb{R}^{d},{\rm vol}_{g_{i}})).$
Hence,
(6.3) $CD\in\mathcal{L}_{d,\infty}(L_{2}(\mathbb{R}^{d},{\rm vol}_{g_{i}})).$
Note that
(6.4) $(1+\Delta_{g_{i}})^{-1}M_{\psi}\in\mathcal{L}_{d,\infty}(L_{2}(\mathbb{R}^{d},{\rm vol}_{g_{i}})).$
Combining (6.1), (6.2), (6.3) and (6.4), we complete the proof. ∎
Lemma 6.6.
For every $\phi\in C^{\infty}(X),$ we have
$M_{\phi},(1+\Delta_{G})^{-1}]\in\mathcal{L}_{\frac{d}{3},\infty}(L_{2}(X,{\rm vol}_{G})).$
Proof.
We have
$[M_{\phi},(1+\Delta_{G})^{-1}]$ $=(1+\Delta_{G})^{-1}[\Delta_{G},M_{\phi}](1+\Delta_{G})^{-1}$
$=(1+\Delta_{G})^{-1}\cdot[\Delta_{G},M_{\phi}](1+\Delta_{G})^{-\frac{1}{2}}\cdot(1+\Delta_{G})^{-\frac{1}{2}}.$
The first factor belongs to $\mathcal{L}_{\frac{d}{2},\infty}(L_{2}(X,{\rm vol}_{G})),$ the second factor is bounded due to Lemma 6.4, the last factor belongs to $\mathcal{L}_{d,\infty}(L_{2}(X,{\rm vol}_{G})).$ The assertion follows from Hölder inequality. ∎
The following theorem is a local version of the Theorem 1.2 with $p=\frac{d}{2}$.
Theorem 6.7.
Let $(X,G)$ be a compact Riemannian manifold and let $\Delta_{G}$ be the Laplace-Beltrami operator. Let $({\mathcal{U}}_{i},h_{i})$ be a chart in a good (in the sense of Definition 6.1) atlas. If $S\in\Pi_{X}$ is compactly supported in ${\mathcal{U}}_{i},$ then
$\lim_{t\rightarrow\infty}t^{\frac{2}{d}}\mu(t,S(1+\Delta_{G})^{-1})=c_{d}\Big\|{\rm sym}_{X}(S)\Big\|_{L_{\frac{d}{2}}(T^{\ast}X,e^{-q_{G}}d\lambda)}$
where $c_{d}$ is given in Theorem A.
Before proving Theorem 6.7, recall a fact that if $H_{1}$ and $H_{2}$ are Hilbert spaces and if $W:H_{1}\to H_{2}$ is a partial isometry, then
$\mu_{B(H_{2})}(WTW^{\ast})=\mu_{B(H_{1})}(T),\quad T\in B(H),\quad T=T(W^{\ast}W)=(W^{\ast}W)T.$
Let us continue to show Theorem 6.7.
Proof of Theorem 6.7.
Fix $\phi\in C^{\infty}_{c}({\mathcal{U}}_{i})$ such that $S=M_{\phi}S=SM_{\phi}.$ Applying the above fact to the operator
$T=S(1+\Delta_{G})^{-1}M_{\phi},$
we conclude
(6.5) $\mu_{B(L_{2}(\mathbb{R}^{d},{\rm vol}_{g_{i}}))}(W_{h_{i}}TW_{h_{i}}^{\ast})=\mu_{B(L_{2}(X,{\rm vol}_{G}))}(T).$
It is immediate that
$W_{h_{i}}TW_{h_{i}}^{\ast}=W_{h_{i}}SW_{h_{i}}^{\ast}\cdot W_{h_{i}}M_{\phi}(1+\Delta_{G})^{-1}M_{\phi}W_{h_{i}}^{\ast}.$
Using Lemma 6.5, we write
$W_{h_{i}}TW_{h_{i}}^{\ast}-W_{h_{i}}SW_{h_{i}}^{\ast}\cdot M_{\phi\circ h_{i}^{-1}}(1+\Delta_{g_{i}})^{-1}M_{\phi\circ h_{i}^{-1}}\in\mathcal{L}_{\frac{d}{3},\infty}(L_{2}(\mathbb{R}^{d},{\rm vol}_{g_{i}})).$
Let $P_{i}=M_{{\rm det}^{-\frac{1}{4}}(g_{i})}\Delta_{g_{i}}M_{{\rm det}^{\frac{1}{4}}(g_{i})}$. By Theorem 3.1,
$M_{{\rm det}^{-\frac{1}{4}}(g_{i})}M_{\phi\circ h_{i}^{-1}}[(1+\Delta_{g_{i}})^{-1},M_{\phi\circ h_{i}^{-1}}]M_{{\rm det}^{\frac{1}{4}}(g_{i})}$ $=M_{\phi\circ h_{i}^{-1}}[(1+P_{i})^{-1},M_{\phi\circ h_{i}^{-1}}]$
$\in\mathcal{L}_{\frac{d}{3},\infty}(L_{2}(\mathbb{R}^{d})),$
and therefore
$M_{\phi\circ h_{i}^{-1}}(1+\Delta_{g_{i}})^{-1}M_{\phi\circ h_{i}^{-1}}-M_{\phi^{2}\circ h_{i}^{-1}}(1+\Delta_{g_{i}})^{-1}\in\mathcal{L}_{\frac{d}{3},\infty}(L_{2}(\mathbb{R}^{d},{\rm vol}_{g_{i}})).$
Since
$W_{h_{i}}SW_{h_{i}}^{\ast}=W_{h_{i}}SW_{h_{i}}^{\ast}\cdot M_{\phi\circ h_{i}^{-1}},$
it follows that
(6.6) $W_{h_{i}}TW_{h_{i}}^{\ast}-W_{h_{i}}SW_{h_{i}}^{\ast}\cdot(1+\Delta_{g_{i}})^{-1}\in\mathcal{L}_{\frac{d}{3},\infty}(L_{2}(\mathbb{R}^{d},{\rm vol}_{g_{i}})).$
By construction of $\Pi_{X},$ we have ${\rm Ext}_{h_{i}({\mathcal{U}}_{i})}(W_{h_{i}}SW_{h_{i}}^{\ast})\in\Pi_{\mathbb{R}^{d}}.$ By Theorem 4.1, there exists a limit
$\lim_{t\to\infty}t^{\frac{2}{d}}\mu_{B(L_{2}(\mathbb{R}^{d},{\rm vol}_{g_{i}}))}(t,W_{h_{i}}SW_{h_{i}}^{\ast}\cdot(1+\Delta_{g_{i}})^{-1})$
$=c_{d}\|{\rm sym}({\rm Ext}_{h_{i}({\mathcal{U}}_{i})}(W_{h_{i}}SW_{h_{i}}^{\ast}))\|_{L_{\frac{d}{2}}(h_{i}({\mathcal{U}}_{i})\times\mathbb{R}^{d},e^{-q_{g_{i}}}dm)}$
$=c_{d}\|{\rm sym}_{X}(S)\|_{L_{\frac{d}{2}}(T^{\ast}X,e^{-q_{G}}d\lambda)}.$
Using Lemma 2.5 and taking (6.6) into account, we conclude that
$\lim_{t\to\infty}t^{\frac{2}{d}}\mu_{B(L_{2}(\mathbb{R}^{d},{\rm vol}_{g_{i}}))}(t,W_{h_{i}}TW_{h_{i}}^{\ast})=c_{p,d}\|{\rm sym}_{X}(S)\|_{L_{p}(T^{\ast}X,e^{-q_{G}}d\lambda)}.$
Using (6.5), we write
(6.7) $\lim_{t\rightarrow\infty}t^{\frac{2}{d}}\mu(t,T)=c_{d}\Big\|{\rm sym}_{X}(S)\Big\|_{L_{\frac{d}{2}}(T^{\ast}X,e^{-q_{G}}d\lambda)}.$
Since $S=SM_{\phi}$ and since (from Lemma 6.6)
$M_{\phi}(1+\Delta_{G})^{-1}M_{\phi}-M_{\phi^{2}}(1+\Delta_{G})^{-1}\in\mathcal{L}_{\frac{d}{3},\infty}(L_{2}(X,{\rm vol}_{G})),$
it follows that
$T-S(1+\Delta_{G})^{-1}\in\mathcal{L}_{\frac{d}{3},\infty}(L_{2}(X,{\rm vol}_{G})).$
The assertion follows now from (6.7) and Lemma 2.5. ∎
7. Asymptotics on manifold
In this section, we complete the proof of Theorem 1.2. The case where $p=\frac{d}{2}$ is treated in Subsection 7.1, while the general case is addressed in Subsection 7.2. In this section, let $c_{d}$ given in Theorem A in the sequel.
7.1. Asymptotics for $p=\frac{d}{2}$
We prove Theorem 1.2 with $p=\frac{d}{2}$ in this subsection. To do that, we demand the following approximation on manifold.
Lemma 7.1.
Let $X$ be a compact Riemannian manifold with a good atlas $\{({\mathcal{U}}_{i},h_{i})\}_{i\in\mathbb{I}}$. There is a sequence $\{\phi_{n,k}\}_{n\in\mathbb{N},1\leq k\leq K}$ such that
  1. (a)
    for all $n\in\mathbb{N}$ and $1\leq k\leq K$, function $\phi_{n,k}$ is a smooth function on $X$;
  2. (b)
    for all $n\in\mathbb{N}$ and $1\leq k\leq K$, function $0\leq\phi_{n,k}\leq 1$;
  3. (c)
    for all $n\in\mathbb{N},1\leq k\leq K$, function $\phi_{n,k}$ is supported in some chart $({\mathcal{U}}_{i_{k}},h_{i_{k}})$;
  4. (d)
    the sum $\sum_{k=1}^{K}\phi_{n,k}^{3}\rightarrow 1$ in the sense of almost everywhere as $n\rightarrow\infty$;
  5. (e)
    for all $n\in\mathbb{N}$, the functions $\{\phi_{n,k}\}_{k=1}^{K}$ are disjoint supported.
Proof.
Choose pairwise disjoint open sets $\{\Delta_{k}\}_{k=1}^{K}$ such that $\Delta_{k}$ is contained in some $\mathcal{U}_{i_{k}}$ and their union has full measure. For $1\leq k\leq K$, choose a sequence $\{\phi_{n,k}\}_{n\in\mathbb{N}}$ such that (i) $\phi_{n,k}\in C^{\infty}_{c}(\Delta_{k})$ is positive taking values in $[0,1]$ for every $n\in\mathbb{N}$ and (ii) $\phi_{n,k}\uparrow\chi_{\Delta_{k}}$ almost everywhere as $n\to\infty.$ The assertions follow. ∎
Lemma 7.2.
Let $\phi\in C^{\infty}(X)$ be supported in some chart $({\mathcal{U}}_{i},h_{i})$ (which belongs to a good atlas). For any $S\in\Pi_{X}$, we have
$\lim_{t\rightarrow\infty}t^{\frac{2}{d}}\mu(t,M_{\phi}SM_{\phi}(1+\Delta_{G})^{-1}M_{\phi})=c_{d}\|{\rm sym}_{X}(M_{\phi^{3}}S)\|_{L_{\frac{d}{2}}(T^{\ast}X,e^{-q_{G}}d\lambda)}.$
Proof.
Applying Theorem 6.7 to the operator $M_{\phi}SM_{\phi^{2}},$ we conclude that
$\lim_{t\rightarrow\infty}t^{\frac{1}{p}}\mu(t,M_{\phi}SM_{\phi^{2}}(1+\Delta_{G})^{-\frac{d}{2p}})=c_{p,d}\Big\|{\rm sym}_{X}(M_{\phi}SM_{\phi^{2}})\Big\|_{L_{p}(T^{\ast}X,e^{-q_{G}}d\lambda)}.$
Write
$M_{\phi}SM_{\phi}(1+\Delta_{G})^{-1}M_{\phi}=M_{\phi}SM_{\phi^{2}}(1+\Delta_{G})^{-1}+M_{\phi}SM_{\phi}\cdot[(1+\Delta_{G})^{-1},M_{\phi}].$
By Lemma 6.6, we have
$M_{\phi}SM_{\phi}[(1+\Delta_{G})^{-1},M_{\phi}]\in\mathcal{L}_{\frac{d}{3},\infty}(L_{2}(X,{\rm vol}_{G})).$
Therefore, by Lemma 2.5,
$\lim_{t\rightarrow\infty}t^{\frac{2}{d}}\mu(t,M_{\phi}SM_{\phi}(1+\Delta_{G})^{-1}M_{\phi})=c_{d}\|{\rm sym}_{X}(M_{\phi}SM_{\phi^{2}})\|_{L_{\frac{d}{2}}(T^{\ast}X,e^{-q_{G}}d\lambda)}.$
Taking into account that
${\rm sym}_{X}(M_{\phi}SM_{\phi^{2}})$ $={\rm sym}_{X}(M_{\phi})\cdot{\rm sym}_{X}(S)\cdot{\rm sym}_{S}(M_{\phi})^{2}$
$={\rm sym}_{X}(M_{\phi^{3}})\cdot{\rm sym}_{X}(S)$
$={\rm sym}_{X}(M_{\phi^{3}}S),$
we complete the proof. ∎
Lemma 7.3.
For each $n\in\mathbb{N}$, let $\psi_{n}=\sum_{k=1}^{K}\phi_{n,k}^{3}$ where sequence $\{\phi_{n,k}\}_{n\in\mathbb{N},1\leq k\leq K}$ is given in Lemma 7.1. For any $S\in\Pi_{X}$, we have
$\lim_{t\rightarrow\infty}t^{\frac{2}{d}}\mu(t,SM_{\psi_{n}}(1+\Delta_{G})^{-1})=c_{d}\|{\rm sym}_{X}(M_{\psi_{n}}S)\|_{L_{\frac{d}{2}}(T^{\ast}X,e^{-q_{G}}d\lambda)}.$
Proof.
By Lemma 7.2, for every $n\in\mathbb{N}$ and for every $1\leq k\leq K,$ there exists a limit
$\lim_{t\rightarrow\infty}t^{\frac{2}{d}}\mu(t,M_{\phi_{n,k}}SM_{\phi_{n,k}}(1+\Delta_{G})^{-1}M_{\phi_{n,k}})=c_{d}\|{\rm sym}_{X}(M_{\phi_{n,k}^{3}}S)\|_{L_{\frac{d}{2}}(T^{\ast}X,e^{-q_{G}}d\lambda)}.$
For every $n\in\mathbb{N},$ the operators
$\{M_{\phi_{n,k}}SM_{\phi_{n,k}}(1+\Delta_{G})^{-1}M_{\phi_{n,k}}\}_{1\leq k\leq K}$
are pairwise orthogonal. By Lemma 2.4, there exists a limit
$\lim_{t\rightarrow\infty}t^{\frac{2}{d}}\mu(t,\sum_{k=1}^{K}M_{\phi_{n,k}}SM_{\phi_{n,k}}(1+\Delta_{G})^{-1}M_{\phi_{n,k}})$
$=c_{d}\Big(\sum_{k=1}^{K}\|{\rm sym}_{X}(M_{\phi_{n,k}^{3}}S)\|_{L_{\frac{d}{2}}(T^{\ast}X,e^{-q_{G}}d\lambda)}^{\frac{d}{2}}\Big)^{\frac{2}{d}}.$
Write
$SM_{\psi_{n}}(1+\Delta_{G})^{-1}$ $=\sum_{k=1}^{K}SM_{\phi_{n,k}^{3}}(1+\Delta_{G})^{-1}$
$=\sum_{k=1}^{K}[S,M_{\phi_{n,k}}]\cdot M_{\phi_{n,k}^{2}}(1+\Delta_{G})^{-1}+$
$\quad+\sum_{k=1}^{K}M_{\phi_{n,k}}SM_{\phi_{n,k}}\cdot[M_{\phi_{n,k}},(1+\Delta_{G})^{-1}]+$
$\quad+\sum_{k=1}^{K}M_{\phi_{n,k}}SM_{\phi_{n,k}}(1+\Delta_{G})^{-1}M_{\phi_{n,k}}.$
Note that $S,M_{\phi_{n,k}}]$ is compact (this is part of the definition of $\Pi_{X}$). Hence,
$S,M_{\phi_{n,k}}]\cdot M_{\phi_{n,k}^{2}}(1+\Delta_{G})^{-1}\in(\mathcal{L}_{\frac{d}{2},\infty})_{0}(L_{2}(X,{\rm vol}_{G})).$
By Lemma 6.6
$M_{\phi_{n,k}}SM_{\phi_{n,k}}\cdot[M_{\phi_{n,k}},(1+\Delta_{G})^{-1}]\in(\mathcal{L}_{\frac{d}{2},\infty})_{0}(L_{2}(X,{\rm vol}_{G})).$
Hence,
$SM_{\psi_{n}}(1+\Delta_{G})^{-1}-\sum_{k=1}^{K}M_{\phi_{n,k}}SM_{\phi_{n,k}}(1+\Delta_{G})^{-1}M_{\phi_{n,k}}\in(\mathcal{L}_{\frac{d}{2},\infty})_{0}(L_{2}(X,{\rm vol}_{G})).$
By Lemma 2.5,
$\lim_{t\rightarrow\infty}t^{\frac{2}{d}}\mu(t,SM_{\psi_{n}}(1+\Delta_{G})^{-1})$
$=c_{d}(\sum_{k=1}^{K}\|{\rm sym}_{X}(M_{\phi_{n,k}^{3}}S)\|_{L_{\frac{d}{2}}(T^{\ast}X,e^{-q_{G}}d\lambda)}^{\frac{d}{2}})^{\frac{2}{d}}$
$=c_{d}(\sum_{k=1}^{K}\|{\rm sym}_{X}(M_{\phi_{n,k}})^{3}\cdot{\rm sym}_{X}(S)\|_{L_{\frac{d}{2}}(T^{\ast}X,e^{-q_{G}}d\lambda)}^{\frac{d}{2}})^{\frac{2}{d}}.$
Since the functions $\{\phi_{n,k}\}_{1\leq k\leq K}$ are pairwise disjointly supported, it follows that so are the functions $\{{\rm sym}_{X}(M_{\phi_{n,k}})\}_{1\leq k\leq K}.$ Hence,
$(\sum_{k=1}^{K}\|{\rm sym}_{X}(M_{\phi_{n,k}})^{3}\cdot{\rm sym}_{X}(S)\|_{L_{\frac{d}{2}}(T^{\ast}X,e^{-q_{G}}d\lambda)}^{\frac{d}{2}})^{\frac{2}{d}}$
$=\|{\rm sym}_{X}(M_{\psi_{n}})\cdot{\rm sym}_{X}(S)\|_{L_{\frac{d}{2}}(T^{\ast}X,e^{-q_{G}}d\lambda)}.$
This completes the proof. ∎
We firstly provide the proof for Theorem 1.2 in the case $p=\frac{d}{2}$.
Proof of Theorem 1.2 for $p=\frac{d}{2}$.
For each $n\in\mathbb{N}$, let $\psi_{n}=\sum_{k=1}^{K}\phi_{n,k}^{3}$ where the sequence $\{\phi_{n,k}\}_{n\in\mathbb{N},1\leq k\leq K}$ is given in Lemma 7.1. It is immediate that
$\|SM_{\psi_{n}}(1+\Delta_{G})^{-1}-S(1+\Delta_{G})^{-1}\|_{\mathcal{L}_{\frac{d}{2},\infty}(L_{2}(X,{\rm vol}_{G}))}$
$\leq\|S\|_{\infty}\|M_{\psi_{n}-1}(1+\Delta_{G})^{-1}\|_{\mathcal{L}_{\frac{d}{2},\infty}(L_{2}(X,{\rm vol}_{G}))}.$
Applying Lemma 2.3, we conclude
$\|M_{\psi_{n}-1}(1+\Delta_{G})^{-1}\|_{\mathcal{L}_{\frac{d}{2},\infty}}\leq c_{X,g}\|\psi_{n}-1\|_{\max\{\frac{d}{2},3\}}\to 0,\quad n\to\infty.$
Here, the convergence is guaranteed by Lemma 7.1 and by the Dominated Convergence Theorem.
Lemma 7.3 asserts that
$\lim_{t\rightarrow\infty}t^{\frac{2}{d}}\mu(t,SM_{\psi_{n}}(1+\Delta_{G})^{-1})=c_{d}\|{\rm sym}_{X}(M_{\psi_{n}}S)\|_{L_{\frac{d}{2}}(T^{\ast}X,e^{-q_{G}}d\lambda)}.$
Using Lemma 2.6, we conclude that the following limits exist and are equal:
$\lim_{t\rightarrow\infty}t^{\frac{2}{d}}\mu(t,S(1+\Delta_{G})^{-1})$ $=c_{d}\lim_{n\to\infty}\|{\rm sym}_{X}(M_{\psi_{n}}S)\|_{L_{\frac{d}{2}}(T^{\ast}X,e^{-q_{G}}d\lambda)}$
$=c_{d}\|{\rm sym}_{X}(S)\|_{L_{\frac{d}{2}}(T^{\ast}X,e^{-q_{G}}d\lambda)}.$
Here, the very last equality is guaranteed by Lemma 7.1 and by the Dominated Convergence Theorem. This completes the proof. ∎
7.2. General case for an arbitrary $p$
In this subsection, we complete the proof of Theorem 1.2.
Lemma 7.4.
Let $P:W^{2,2}(\mathbb{R}^{d})\to L_{2}(\mathbb{R}^{d})$ be a positive, self-adjoint, elliptic differential operator whose coefficients are smooth and also are constant outside some ball. For every $\psi\in C_{0}(\mathbb{R}^{d})$ and for every $\phi\in C^{\infty}_{c}(\mathbb{R}^{d})$ we have
$M_{\psi},M_{\phi}(1+P)^{-1}M_{\phi}]\in(\mathcal{L}_{\frac{d}{2},\infty})_{0}(L_{2}(\mathbb{R}^{d})).$
Proof.
Write
$[M_{\psi},M_{\phi}(1+P)^{-1}M_{\phi}]=[M_{\psi\phi},(1+P)^{-1}]M_{\phi}+[(1+P)^{-1},M_{\phi}]M_{\psi\phi}.$
By theorem 3.1, the factors on the right hand side belong to $\mathcal{L}_{\frac{d}{3},\infty}(L_{2}(\mathbb{R}^{d}))$ and therefore $(\mathcal{L}_{\frac{d}{2},\infty})_{0}(L_{2}(\mathbb{R}^{d}))$. ∎
Lemma 7.5.
Let $P:W^{2,2}(\mathbb{R}^{d})\to L_{2}(\mathbb{R}^{d})$ be a positive, self-adjoint, elliptic differential operator whose coefficients are smooth and also are constant outside some ball. For every $\phi\in C^{\infty}_{c}(\mathbb{R}^{d})$ we have
$R_{j},M_{\phi}(1+P)^{-1}M_{\phi}]\in(\mathcal{L}_{\frac{d}{2},\infty})_{0}(L_{2}(\mathbb{R}^{d})),\quad 1\leq j\leq d.$
Proof.
Write
$[R_{j},M_{\phi}(1+P)^{-1}M_{\phi}]=R_{j}M_{\phi}[(1+P)^{-1},M_{\phi}]+[(1+P)^{-1},M_{\phi}]M_{\phi}R_{j}+$
$+[R_{j},M_{\phi}]M_{\phi}(1+P)^{-1}+(1+P)^{-1}M_{\phi}[R_{j},M_{\phi}]+$
$+[M_{\phi}R_{j}M_{\phi},(1+P)^{-1}].$
By Theorem 3.1 the first and second factors on the right hand side belong to $\mathcal{L}_{\frac{d}{3},\infty}(L_{2}(\mathbb{R}^{d}))$ and therefore $(\mathcal{L}_{\frac{d}{2},\infty})_{0}(L_{2}(\mathbb{R}^{d}))$. Since operator $R_{j},M_{\phi}]$ is compact (see e.g. [41, 20]) and $M_{\phi}(1+P)^{-1},(1+P)^{-1}M_{\phi}$ belong to $\mathcal{L}_{\frac{d}{2},\infty}(L_{2}(\mathbb{R}^{d}))$ by Lemma 3.11, the third and fourth factors on the right hand side belong to $(\mathcal{L}_{\frac{d}{2},\infty})_{0}(L_{2}(\mathbb{R}^{d}))$. For the fifth factor on the right hand side, write
$[M_{\phi}R_{j}M_{\phi},(1+P)^{-1}]=$
$=(1+P)^{-\frac{1}{2}}[M_{\phi}R_{j}M_{\phi},(1+P)^{-\frac{1}{2}}]+[M_{\phi}R_{j}M_{\phi},(1+P)^{-\frac{1}{2}}](1+P)^{-\frac{1}{2}}.$
Applying Theorem 3.2, it gives that operator $M_{\phi}R_{j}M_{\phi},(1+P)^{-1}]$ belongs to $\mathcal{L}_{\frac{d}{3},\infty}(L_{2}(\mathbb{R}^{d}))$ and therefore $(\mathcal{L}_{\frac{d}{2},\infty})_{0}(L_{2}(\mathbb{R}^{d}))$. The proof is complete. ∎
Lemma 7.6.
Let $P:W^{2,2}(\mathbb{R}^{d})\to L_{2}(\mathbb{R}^{d})$ be a positive, self-adjoint, elliptic differential operator whose coefficients are smooth and also are constant outside some ball. For every $S\in\Pi_{\mathbb{R}^{d}}$ and for every $\phi\in C^{\infty}_{c}(\mathbb{R}^{d})$ we have
$S,M_{\phi}(1+P)^{-1}M_{\phi}]\in(\mathcal{L}_{\frac{d}{2},\infty})_{0}(L_{2}(\mathbb{R}^{d})).$
Proof.
Let $S$ belong to $\ast$-algebra generated by $\{M_{f}:\ f\in C_{0}(\mathbb{R}^{d})\}$ and by $\{R_{j}\}_{j=1}^{d}.$ Using Lemmas 7.4, 7.5 and the Leibniz rule, we conclude that
$S,M_{\phi}(1+P)^{-1}M_{\phi}]\in(\mathcal{L}_{\frac{d}{2},\infty})_{0}(L_{2}(\mathbb{R}^{d})).$
Consider now the general case. By the definition of $\Pi_{\mathbb{R}^{d}},$ we can find a sequence $\{S_{n}\}_{n\geq 0}$ such that each $S_{n}$ belongs to $\ast$-algebra generated by $\{M_{f}:\ f\in C_{0}(\mathbb{R}^{d})\}$ and $\{R_{j}\}_{j=1}^{d}$, and such that $S_{n}\to S$ in the uniform norm. By the preceding paragraph,
$S_{n},M_{\phi}(1+P)^{-1}M_{\phi}]\in(\mathcal{L}_{\frac{d}{2},\infty})_{0}(L_{2}(\mathbb{R}^{d})),\quad n\geq 0.$
We clearly have
$S_{n},M_{\phi}(1+P)^{-1}M_{\phi}]\to[S,M_{\phi}(1+P)^{-1}M_{\phi}],\quad n\to\infty.$
Since $(\mathcal{L}_{\frac{d}{2},\infty})_{0}(L_{2}(\mathbb{R}^{d}))$ is closed in $\mathcal{L}_{\frac{d}{2},\infty}(L_{2}(\mathbb{R}^{d})),$ the assertion follows. ∎
Lemma 7.7.
Let $g:\mathbb{R}^{d}\to{\rm GL}_{+}(d,\mathbb{R})$ be smooth mapping constant outside some ball. For every $S\in\Pi_{\mathbb{R}^{d}}$ and for every $\phi\in C^{\infty}_{c}(\mathbb{R}^{d}),$ we have
$S,M_{\phi}(1+\Delta_{g})^{-1}M_{\phi}]\in(\mathcal{L}_{\frac{d}{2},\infty})_{0}(L_{2}(\mathbb{R}^{d},{\rm vol}_{g})).$
Proof.
Consider the unitary mapping $M_{{\rm det}^{\frac{1}{4}}(g)}:L_{2}(\mathbb{R}^{d},{\rm vol}_{g})\to L_{2}(\mathbb{R}^{d}).$ We have
$M_{{\rm det}^{-\frac{1}{4}}(g)}[S,M_{\phi}(1+\Delta_{g})^{-1}M_{\phi}]M_{{\rm det}^{\frac{1}{4}}(g)}=[M_{{\rm det}^{-\frac{1}{4}}(g)}SM_{{\rm det}^{\frac{1}{4}}(g)},M_{\phi}(1+P)^{-1}M_{\phi}]$
where
$P=M_{{\rm det}^{-\frac{1}{4}}(g)}\Delta_{g}M_{{\rm det}^{\frac{1}{4}}(g)}.$
Since $P:W^{2,2}(\mathbb{R}^{d})\to L_{2}(\mathbb{R}^{d})$ be a positive, self-adjoint, elliptic differential operator whose coefficients are smooth and also are constant outside some ball, it follows from Lemma 7.6 that
$M_{{\rm det}^{-\frac{1}{4}}(g)}SM_{{\rm det}^{\frac{1}{4}}(g)},M_{\phi}(1+P)^{-1}M_{\phi}]\in(\mathcal{L}_{\frac{d}{2},\infty})_{0}(L_{2}(\mathbb{R}^{d})).$
This suffices to complete the proof. ∎
Lemma 7.8.
Let $(X,G)$ be a compact Riemannian manifold and let $\Delta_{G}$ be the Laplace-Beltrami operator. Let $({\mathcal{U}}_{i},h_{i})$ be a chart in a good atlas. For every $S\in\Pi_{X}$ compactly supported in ${\mathcal{U}}_{i},$ we have
$S,(1+\Delta_{G})^{-1}]\in(\mathcal{L}_{\frac{d}{2},\infty})_{0}(L_{2}(X,{\rm vol}_{G})).$
Proof.
Fix $\phi\in C^{\infty}_{c}({\mathcal{U}}_{i})$ such that $S=M_{\phi}S=SM_{\phi}.$ We have
$[S,(1+\Delta_{G})^{-1}]$ $=S(1+\Delta_{G})^{-1}-(1+\Delta_{G})^{-1}S$
$=SM_{\phi^{2}}(1+\Delta_{G})^{-1}-(1+\Delta_{G})^{-1}M_{\phi^{2}}S$
$=[S,M_{\phi}(1+\Delta_{G})^{-1}M_{\phi}]+S\cdot[M_{\phi},(1+\Delta_{G})^{-1}]+$
$\quad+[M_{\phi},(1+\Delta_{G})^{-1}]\cdot S.$
By Lemma 6.6,
$M_{\phi},(1+\Delta_{G})^{-1}]\in\mathcal{L}_{\frac{d}{3},\infty}(L_{2}(X,{\rm vol}_{G})).$
Thus,
$S,(1+\Delta_{G})^{-1}]-[S,M_{\phi}(1+\Delta_{G})^{-1}M_{\phi}]\in\mathcal{L}_{\frac{d}{3},\infty}(L_{2}(X,{\rm vol}_{G})).$
Lemma 6.5 asserts that
$M_{\phi}(1+\Delta_{G})^{-1}M_{\phi}-W_{h_{i}}^{\ast}M_{\phi\circ h_{i}^{-1}}(1+\Delta_{g_{i}})^{-1}M_{\phi\circ h_{i}^{-1}}W_{h_{i}}\in\mathcal{L}_{\frac{d}{3},\infty}(L_{2}(X,{\rm vol}_{G})).$
Thus,
$S,(1+\Delta_{G})^{-1}]-[S,W_{h_{i}}^{\ast}M_{\phi\circ h_{i}^{-1}}(1+\Delta_{g_{i}})^{-1}M_{\phi\circ h_{i}^{-1}}W_{h_{i}}]\in\mathcal{L}_{\frac{d}{3},\infty}(L_{2}(X,{\rm vol}_{G})).$
In other words,
$S,(1+\Delta_{G})^{-1}]-W_{h_{i}}^{\ast}[S_{i},M_{\phi\circ h_{i}^{-1}}(1+\Delta_{g_{i}})^{-1}M_{\phi\circ h_{i}^{-1}}]W_{h_{i}}\in\mathcal{L}_{\frac{d}{3},\infty}(L_{2}(X,{\rm vol}_{G})),$
where $S_{i}=W_{h_{i}}SW_{h_{i}}^{\ast}\in\Pi_{\mathbb{R}^{d}}.$ The assertion follows now from Lemma 7.7. ∎
Lemma 7.9.
Let $(X,G)$ be a compact Riemannian manifold and let $\Delta_{G}$ be the Laplace-Beltrami operator. For every $S\in\Pi_{X},$ we have
$S,(1+\Delta_{G})^{-1}]\in(\mathcal{L}_{\frac{d}{2},\infty})_{0}(L_{2}(X,{\rm vol}_{G})).$
Proof.
Let $\{(U_{n},h_{n})\}_{n=1}^{N}$ be a good atlas and let $\{\phi_{k}\}_{k=1}^{K}$ be a partition of unity such that, for every $1\leq k_{1},k_{2}\leq K,$ either there exists $n_{k_{1},k_{2}}$ such that $\phi_{k_{1}},\phi_{k_{2}}\in C_{c}({\mathcal{U}}_{n_{k_{1},k_{2}}})$ or $\phi_{k_{1}}\phi_{k_{2}}\equiv 0.$ We write
$S=\sum_{k_{1},k_{2}=1}^{K}M_{\phi_{k_{1}}}SM_{\phi_{k_{2}}}.$
If there exists $n_{k_{1},k_{2}}$ such that $\phi_{k_{1}},\phi_{k_{2}}\in C^{\infty}_{c}({\mathcal{U}}_{n_{k_{1},k_{2}}}),$ then
$M_{\phi_{k_{1}}}SM_{\phi_{k_{2}}},(1+\Delta_{G})^{-1}]\in(\mathcal{L}_{\frac{d}{2},\infty})_{0}(L_{2}(X,{\rm vol}_{G}))$
by Lemma 7.8. If $\phi_{k_{1}}\phi_{k_{2}}=0,$ then
$M_{\phi_{k_{1}}}SM_{\phi_{k_{2}}}=M_{\phi_{k_{1}}}\cdot[S,M_{\phi_{k_{2}}}]\in\mathcal{K}(L_{2}(X,{\rm vol}_{G})).$
In either case, we have
$M_{\phi_{k_{1}}}SM_{\phi_{k_{2}}},(1+\Delta_{G})^{-1}]\in(\mathcal{L}_{\frac{d}{2},\infty})_{0}(L_{2}(X,{\rm vol}_{G})).$
Summing over $1\leq k_{1},k_{2}\leq K,$ we complete the proof. ∎
Now, let us give a complete proof for Theorem 1.2.
Proof of Theorem 1.2.
Without loss of generality, $S\geq 0.$ Using Theorem 1.2 for the case $p=\frac{d}{2},$ we write
$\lim_{t\rightarrow\infty}t^{\frac{2}{d}}\mu(t,S^{\frac{2p}{d}}(1+\Delta_{G})^{-1})=c_{d}\|{\rm sym}_{X}(S^{\frac{2p}{d}})\|_{L_{\frac{d}{2}}(T^{\ast}X,e^{-q_{G}}d\lambda)}.$
According to Lemma 7.9,
$S^{\frac{2p}{d}}(1+\Delta_{G})^{-1}-S^{\frac{p}{d}}(1+\Delta_{G})^{-1}S^{\frac{p}{d}}\in(\mathcal{L}_{\frac{d}{2},\infty})_{0}(L_{2}(X,{\rm vol}_{G})).$
It follows from Lemma 2.5 that
$\lim_{t\rightarrow\infty}t^{\frac{2}{d}}\mu(t,S^{\frac{p}{d}}(1+\Delta_{G})^{-1}S^{\frac{p}{d}})=c_{d}\|{\rm sym}_{X}(S^{\frac{2p}{d}})\|_{L_{\frac{d}{2}}(T^{\ast}X,e^{-q_{G}}d\lambda)}.$
Taking this equality to the power $\frac{d}{2p},$ we write
$\lim_{t\rightarrow\infty}t^{\frac{1}{p}}\mu\Big(t,\Big(S^{\frac{p}{d}}(1+\Delta_{G})^{-1}S^{\frac{p}{d}}\Big)^{\frac{d}{2p}}\Big)=c_{d}^{\frac{d}{2p}}\|{\rm sym}_{X}(S^{\frac{2p}{d}})\|_{L_{\frac{d}{2}}(T^{\ast}X,e^{-q_{G}}d\lambda)}^{\frac{d}{2p}}.$
Applying Theorem 5.1 with $s=\frac{d}{2p}$ and with $S^{\frac{2p}{d}}$ instead of $S,$ we write
$S(1+\Delta_{G})^{-\frac{d}{2p}}-(S^{\frac{p}{d}}(1+\Delta_{G})^{-1}S^{\frac{p}{d}})^{\frac{d}{2p}}\in(\mathcal{L}_{p,\infty})_{0}(L_{2}(X,{\rm vol}_{G})).$
It follows from Lemma 2.5 that
$\lim_{t\rightarrow\infty}t^{\frac{1}{p}}\mu\Big(t,S(1+\Delta_{G})^{-\frac{d}{2p}}\Big)=c_{d}^{\frac{d}{2p}}\|{\rm sym}_{X}(S^{\frac{2p}{d}})\|_{L_{\frac{d}{2}}(T^{\ast}X,e^{-q_{G}}d\lambda)}^{\frac{d}{2p}}.$
Since ${\rm sym}_{X}$ is an $\ast$-homomorphism, it follows that
${\rm sym}_{X}(S^{\frac{2p}{d}})={\rm sym}_{X}(S)^{\frac{2p}{d}}$
and the assertion follows. The proof is complete. ∎
Acknowledgement: The first and the third author are supported by the ARC DP230100434.